### v1

# CrochetBench: Can Vision-Language Models Move from Describing to Doing in Crochet Domain?

Abstract
We present CrochetBench, a benchmark for evaluating the ability of multimodal large language models to perform fine-grained, low-level procedural reasoning in the domain of crochet. Unlike prior benchmarks that focus on high-level description or visual question answering, CrochetBench shifts the emphasis from describing to doing: models are required to recognize stitches, select structurally appropriate instructions, and generate compilable crochet procedures. We adopt the CrochetPARADE DSL as our intermediate representation, enabling structural validation and functional evaluation via execution. The benchmark covers tasks including stitch classification, instruction grounding, and both natural language and image-to-DSL translation. Across all tasks, performance sharply declines as the evaluation shifts from surface-level similarity to executable correctness, exposing limitations in long-range symbolic reasoning and 3D-aware procedural synthesis. CrochetBench offers a new lens for assessing procedural competence in multimodal models and highlights the gap between surface-level understanding and executable precision in real-world creative domains. Code is available at https://github.com/Peiyu-Georgia-Li/crochetBench.
1 Introduction
Procedural crafts such as crochet present a distinctive frontier for multimodal learning. Unlike traditional captioning or recipe datasets (chefFusion; Hu_2023_ICCV; mohbat2024llava), crochet patterns intertwine three interdependent modalities: (i) structured symbolic language, where stitch abbreviations and counts define a precise grammar of construction; (ii) long-form natural language, which provides contextual guidance such as materials and sizing; and (iii) visual evidence, including photographs of completed objects and motif diagrams. Success requires not just alignment across modalities but step-wise reasoning that preserves procedural fidelity, making the challenge closer to program synthesis than generic description.
Crochet also offers a unique testbed for 3D-aware reasoning. Each stitch encodes both local geometry and global connectivity, forming a topological structure that must be preserved across steps. Generating or interpreting patterns thus demands reasoning over how sequential operations accumulate into volumetric form. In effect, crochet couples symbolic instruction following with embodied spatial reasoning, cultivating abilities essential for domains where language must ground into physical tasks.
Despite the rapid growth of multimodal benchmarks (fu2024mmecomprehensiveevaluationbenchmark; li2023seed; zhang-etal-2025-ngqa; yue2024mmmu), existing datasets have largely focused on description or grounding. COCO (lin2015microsoftcococommonobjects) catalyzed captioning research, TextCaps (sidorov2020textcapsdatasetimagecaptioning) extended it to text-in-the-wild, and Recipe1M (marin2019recipe1mdatasetlearningcrossmodal) explored cross-modal cooking instructions. While recipes also involve multi-step procedures, validating correctness typically requires real-world execution, making large-scale evaluation slow and resource-intensive. Crochet, by contrast, provides a symbolic domain where outputs can be automatically verified through DSL compilation, enabling scalable and efficient study of step-wise reasoning. For a more detailed survey, see Appendix A. Yet these benchmarks stop short of testing whether models can follow symbolic grammars, respect numerical and spatial constraints, and produce outputs that are executable. Current systems can describe, but not reliably do.
CrochetBench fills this gap by centering evaluation on instructional fidelity: can models not only recognize and generate, but also output step-wise, compilable instructions that respect symbolic, numerical, and topological structure? Each example in CrochetBench is a multimodal package—structured JSON metadata (stitch inventories and abbreviations), full-text procedures with rows/rounds and conditionals, and paired images of finished objects and motifs. Crucially, CrochetBench is paired with CrochetPARADE (crochetparade), a domain-specific language (DSL) enabling executable evaluation, where natural language instructions are translated into compilable code enforcing geometric and topological coherence.
Our contributions are fourfold: (1) CrochetBench, the first executable benchmark for procedural crafts, unifying symbolic, textual, and visual modalities with evaluation protocols emphasizing procedural fidelity and 3D-aware reasoning; (2) a comprehensive task suite spanning recognition, comprehension, generation, and DSL translation; (3) integration of CrochetPARADE into an executable pipeline, enabling scalable, automated verification of outputs—unlike domains such as cooking, which require real-world execution—thereby shifting evaluation from surface similarity to procedural fidelity; and (4) baseline analyses of state-of-the-art VLMs/MLLMs, revealing systematic weaknesses including hallucinations, captioning bias, and structural artifacts.
Taken together, CrochetBench opens a new direction for multimodal research: moving beyond describing what we see, toward generating executable procedures that respect symbolic grammar, numerical accuracy, and topological coherence—paving the way for models that can reason in structured 3D spaces.
2 Dataset Description
Figure 1: Distribution of the top-10 most common project types in CrochetBench.
CrochetBench is a large-scale, structured benchmark consisting of 6,085 crochet patterns spanning 55 distinct project categories. It was constructed by collecting publicly available patterns from the Yarnspirations website [^1] https://www.yarnspirations.com/collections/patterns, a widely used repository for fiber arts. The raw patterns were originally formatted as PDF documents, which were parsed and normalized using a GPT-4o-mini-based conversion pipeline. This process extracted and standardized key fields including pattern metadata, materials, measurements, gauge, abbreviations, and full step-by-step instructions. Each entry was converted into a machine-readable JSON object with a consistent schema, and 98.77% of patterns include an associated product image, enabling multimodal supervision.
Figure 2: Skill level distribution across the CrochetBench dataset. Note that the “Experienced” slice (1.3%) is annotated externally due to its small size.
The dataset supports diverse real-world crochet practices, with project types ranging from simple accessories to complex garments. Figure 1 lists the ten most common categories by frequency. The majority of patterns belong to a small number of dominant types—Afghans and Blankets alone account for over one-quarter of the dataset. More details can be found at Appendix B.
Each pattern is labeled with one of four primary skill levels—beginner, easy, intermediate, or experienced. This allows for stratified evaluation across complexity tiers. Figure 2 shows the skill level distribution, which is strongly skewed toward beginner-friendly content. Only one pattern (0.02%) is missing a skill level label. More details can be found at Appendix B.
Instructional complexity varies substantially across patterns. The number of characters in each instruction ranges from 20 to over 30,000, with a mean of 3,216 and a median of 2,453. Abbreviation counts (i.e., unique stitch tokens per pattern) range from 1 to 31, with an average of 10.6. These statistics are summarized in Appendix B.1. We observe a clear correlation between skill level and instruction length: beginner patterns tend to be short and use fewer abbreviations, while experienced patterns are significantly longer and more symbolically dense.
In addition to symbolic complexity, the dataset contains 3,143 abbreviation instances mapped to 789 unique standardized stitch tokens. This lexical mapping enables tasks such as vocabulary translation, sequence generation, and instruction validation. Beyond raw instructions, the structured schema also records rich metadata, including gauge, hook size, yarn weight, and measurements. A representative dataset entry is provided in Appendix 10.
Overall, CrochetBench provides a rich resource for multimodal modeling, symbolic reasoning, and structure-aware generation. Its coverage across diverse categories and complexity levels enables broad benchmarking of both open-ended generation and instruction fidelity tasks.
Table 1: Overall statistics of the CrochetBench dataset.
Total Patterns Image Coverage Avg. Instr. Length #Project Types
CrochetBench 6,085 98.77% 3,216 characters 55
3 Experiment
We empirically evaluate CrochetBench by defining a set of structured tasks and benchmarking a diverse pool of multimodal large language models. Our experiments are designed to test models across progressively challenging stages, from low-level recognition to high-level executable synthesis.
Benchmark Tasks.
Table 2 summarizes the four evaluation tasks in CrochetBench, which progress systematically from recognition to comprehension, generation, and ultimately executable synthesis. Task A (Stitch Recognition) evaluates a model’s ability to detect symbolic primitives in crochet images, establishing the foundation for multimodal perception. Task B (Instruction Selection) requires models to align visual evidence with candidate textual instructions, thereby testing multimodal grounding and fine-grained comprehension. Unlike conventional description tasks, the candidates are procedural steps rather than captions, and correct selection often requires reasoning about how local steps contribute to the final product. Task C (Instruction Generation) advances from comprehension to open-ended production, challenging models to generate natural-language procedural instructions that are both perceptually grounded and linguistically faithful to domain conventions. Evaluation here emphasizes lexical and symbolic fidelity, but does not directly capture structural validity. Task D (Instruction-to-DSL Translation) addresses this gap by requiring models to output a compilable program in a domain-specific language (DSL). The step-level variant tests local semantic grounding, while the project-level variant demands global structural consistency across the entire pattern. Compilation-based evaluation directly measures executable faithfulness, ensuring that generated instructions are not only linguistically plausible but also structurally sound. Together, these tasks form a structured evaluation ladder that moves from perceptual recognition to programmatic execution, probing both low-level perception and high-level symbolic reasoning.
Table 2: Overview of benchmark tasks in CrochetBench. Tasks progress from recognition to comprehension, generation, and executable synthesis.
ID Ability Tested Task Evaluation Metrics Test Size
A Recognition Stitch Recognition F1, Precision, Recall
6,009
(CrochetBench-A)
B Comprehension Instruction Selection Accuracy
6,003
(CrochetBench-B)
C Generation Instruction Generation BLEU, ROUGE, ChrF
6,009
(CrochetBench-C)
D Formalization Instr.-to-DSL (Step) Valid Pattern Rate
119
(CrochetBench-Dstep)
Instr.-to-DSL (Project) Valid Pattern Rate
100
(CrochetBench-Dproj)
Model Selection.
We evaluate a diverse set of vision–language models (VLMs) spanning both open-source and closed-source families. The open-source group includes Salesforce BLIP-2 Flan-T5 XL (3B), a perception-focused baseline widely used in image–text tasks; Google Gemma 3 (4B) and Qwen2-VL (7B), two recent models trained with large-scale multimodal alignment; and DeepSeek-VL (7B), a larger open-source model designed for stronger vision–language reasoning. On the closed-source side, we assess GPT-4o, Gemini 2.5 Flash-Lite, and Claude Sonnet 4, which represent state-of-the-art commercial VLMs that push the limits of multimodal reasoning. This selection spans perception-heavy baselines to general-purpose multimodal reasoning systems, enabling us to evaluate capabilities across both research-grade and production-grade settings. [^2] We use the term “VLM” broadly to include both traditional vision–language models (e.g., BLIP-2) and modern multimodal large language models (e.g., GPT-4o, Gemini, Claude).
3.1 Task A: Stitch Recognition
The first task evaluates a model’s ability to identify crochet stitch types from an image of a finished product. We construct CrochetBench-A, a subset of 6,009 examples from the full benchmark, where each product image is paired with ground-truth stitch annotations. These labels are derived from the official pattern instructions and normalized into a standardized set of stitch abbreviations (e.g., sc, hdc, dc) to ensure consistency across patterns. Unlike standard image classification, this is a multi-label prediction problem: multiple stitches may co-occur within the same image, often with subtle visual differences in texture and geometry. This task therefore probes fine-grained visual grounding of structured crochet semantics.
Evaluation.
For each example, we compute overlap between the predicted and reference stitch sets. True Positives (TP) are stitches correctly predicted; False Positives (FP) are stitches predicted but not in the reference; and False Negatives (FN) are stitches in the reference but missed by the model. From these counts, we compute precision (fraction of correct predictions among all predictions), recall (fraction of ground-truth stitches recovered), and F1 score (harmonic mean). Metrics are averaged across examples to provide overall performance. This formulation rewards models that recover all present stitches while avoiding spurious predictions.
Accurate stitch recognition is foundational for the benchmark, as later tasks (e.g., instruction selection and instruction generation) depend on robust detection of stitch primitives.
Table 3: Evaluation results on the Stitch Recognition task. We report Precision, Recall, and F1. Best results are bold; second-best are underlined.
Model Size Precision Recall F1
Open Source Salesforce BLIP-2 Flan-T5 XL 3B 0.2953 0.2303 0.2250
Google Gemma 3 4B 0.2054 0.1021 0.1265
DeepSeek-VL 7B 0.5447 0.7476 0.6060
Qwen2-VL 7B 0.5414 0.6974 0.5816
Closed Source GPT-4o 0.6214 0.5939 0.5801
Gemini 2.5 Flash-Lite 0.7449 0.4977 0.5683
Claude Sonnet 4 0.7861 0.5312 0.6094
Results. Claude Sonnet 4 achieves the best overall F1 score, demonstrating strong precision in stitch recognition. Among open-source models, DeepSeek-VL performs best, with notably high recall, while Qwen2-VL offers a competitive balance. These results highlight the gap between commercial VLMs and open-source alternatives, but also suggest that large open-source models are beginning to approach closed-source performance in fine-grained recognition.
3.2 Task B: Instruction Selection
The second task evaluates whether a model can correctly align a finished crochet product image with its corresponding natural-language instruction. We construct CrochetBench-B, a subset of 6,003 examples, where each item consists of one correct instruction and three distractor instructions sampled from the same project category (e.g., hats, rugs). Distractors are carefully chosen to share structural motifs and vocabulary, thereby increasing difficulty and requiring fine-grained visual–textual alignment rather than reliance on superficial cues. The answer distribution across options is approximately uniform (A: 24.9%, B: 25.7%, C: 23.7%, D: 25.7%).
Evaluation.
To enable scalable assessment, we formulate the task as a 4-way multiple-choice question (MCQ). The model must select one option (A–D), with exactly one correct answer. Predictions are extracted using a lightweight regex-based method that identifies explicit letter outputs (e.g., “A”, “The answer is B”). If no parsable choice is found, the response is marked as unanswered. Accuracy is reported as the evaluation metric.
This task provides a controlled measure of visual grounding and semantic alignment between image content and procedural text, without requiring generative modeling. It highlights the challenge of distinguishing between subtle visual cues and domain-specific terminology, which is essential for bridging perception and structured instruction understanding.
Table 4: Evaluation results on the Instruction Selection task (4-way multiple choice). We report Accuracy. Best results are bold; second-best are underlined.
Model Size Accuracy
Open Source Salesforce BLIP-2 Flan-T5 XL 3B 0.2562
Google Gemma 3 4B 0.2494
DeepSeek-VL 7B 0.2892
Qwen2-VL 7B 0.4196
Closed Source GPT-4o 0.5811
Gemini 2.5 Flash-Lite 0.5563
Claude Sonnet 4 0.5739
Results. GPT-4o achieves the highest accuracy, closely followed by Claude and Gemini, all of which substantially outperform open-source VLMs. Among open-source models, Qwen2-VL is the strongest, while BLIP-2 and Gemma remain near chance level, underscoring the difficulty of visually grounded instruction matching in this domain.
3.3 Task C: Instruction Generation
The third task evaluates a model’s ability to generate natural-language crochet instructions directly from an image of a finished product. We construct CrochetBench-C, a subset of 6,009 examples, where each product image is paired with its full pattern. Unlike classification or selection, this task requires producing a multi-step sequence that follows domain-specific syntax, ordering, and stitch logic. Outputs are expected to resemble real-world crochet patterns, written line by line (e.g., “Rnd 1: ch 4, 6 sc in ring”). This setting challenges models to translate visual evidence into coherent procedural text that preserves both symbolic accuracy and structural consistency.
Evaluation.
We evaluate generation quality using standard text-generation metrics and a domain-specific structural measure. BLEU and ROUGE-L capture lexical and n-gram overlap with the reference instructions. ChrF, computed over character n-grams, provides a finer-grained signal of similarity and is particularly suited to stitch abbreviations. While these metrics measure surface-level overlap, they do not directly test whether the generated instructions are structurally consistent (e.g., balanced stitch counts, valid round progression). To address this limitation, we complement Task C with Task D, where compilation-based evaluation of DSL translations directly assesses executable consistency. Together, these tasks allow us to probe both linguistic fidelity and procedural correctness.
Table 5: Evaluation results on the Instruction Generation task. We report BLEU, ROUGE-L, and ChrF. Higher is better. Best results are bold; second-best are underlined.
Model Size BLEU ROUGE-L ChrF
Open Source Salesforce BLIP-2 Flan-T5 XL 3B 0.0021 0.0926 9.32
Google Gemma 3 4B 0.0010 0.0329 5.17
Qwen2-VL 7B 0.0160 0.2084 15.76
DeepSeek-VL 7B 0.0133 0.1968 18.12
Closed Source GPT-4o 0.0333 0.2353 23.80
Gemini 2.5 Flash-Lite 0.0482 0.2583 30.20
Claude Sonnet 4 0.0331 0.2516 22.95
Results. Gemini achieves the strongest overall performance across all metrics, substantially outperforming other closed-source models. GPT-4o and Claude follow closely, while open-source models lag behind with significantly lower BLEU and ChrF scores. This gap highlights the difficulty of generating structurally faithful crochet instructions, which requires models to capture both visual details and domain-specific procedural logic.
3.4 Task D: Instruction-to-DSL Translation
We construct two subsets for Task D: CrochetBench-Dstep (119 items) and CrochetBench-Dproj (100 items). For CrochetBench-Dstep, we manually annotate the correct CrochetPARADE DSL for the previous context of each pattern (a prefix of NL (natural language) –DSL pairs). The model is then given this prefix along with the next natural-language instruction, and must predict the corresponding DSL line. Predictions are evaluated by checking compilation validity with the CrochetPARADE validator. For CrochetBench-Dproj, we adopt a few-shot prompting setup rather than full manual annotation. Annotators provide a single reference program in CrochetPARADE as context, which is paired with a new natural-language instruction and its product image. The model must then generate a complete DSL program, and outputs are assessed by whether they compile fully or partially using the validator.
Design Rationale. Crochet patterns admit many valid DSL realizations (e.g., alternative groupings or equivalent constructs), meaning that no single gold reference is canonical. A reference-based metric could unfairly penalize models that produce semantically correct but structurally different programs. By instead relying on the validator, CrochetBench shifts evaluation away from surface-level string matching toward functional executability—the central criterion for crochet synthesis. This design is consistent with other program synthesis benchmarks, such as semantic parsing and SQL generation, where execution accuracy is often preferred over exact string match.
3.4.1 Step-Level Translation
In the step-level setting, the model receives a prefix of NL–DSL pairs and must generate the DSL line corresponding to the next natural-language instruction. This setup reflects an incremental synthesis process in which correctness depends on maintaining stitch-level consistency across steps. Since crochet patterns are inherently stateful, earlier context is critical for resolving constructs such as increases, repeats, and turning chains. To capture progression through a pattern, we sample 52 early examples (steps 1–2), 34 mid examples (steps 3–4), and 33 late examples (steps 5–6). Step-level inputs are formatted as:
$\text{Prefix (NL--DSL pairs)}+\text{Next NL instruction}\;\;\rightarrow\;\;\text{Next DSL line}.$
This formulation allows models to generate locally plausible DSL steps, but global correctness ultimately requires stronger contextual reasoning.
Evaluation.
We evaluate models using Compilation Success Rate (CSR), defined as the proportion of generated DSL outputs that compile successfully with the CrochetPARADE validator. This validator-based metric is a key strength of CrochetBench: unlike reference-based string matching, CSR directly measures whether generated programs are executable, rewarding functional correctness rather than surface similarity. To better understand systematic errors, we also conduct fine-grained error analysis across four categories: (1) syntax structure errors, (2) stitch definition errors, (3) labeling and reference errors, and (4) structural or formatting issues. A detailed taxonomy with illustrative examples is provided in the Appendix.
CrochetBench further supports multimodal verification, as CrochetPARADE programs can be rendered into simulated product images. These renderings can be compared with ground-truth product images using pretrained vision–language models (e.g., CLIP), enabling evaluation of both structural validity and visual faithfulness. While current model performance leaves ample room for progress, the combination of validator-based execution checks and prospective image-grounded verification establishes CrochetBench as a uniquely rigorous testbed for structured multimodal reasoning.
Table 6: Verification results across models, grouped by category.
Model CSR (%) Undef. Br. MRef Other Tot. Err.
Open Source
Salesforce BLIP-2 Flan-T5 XL 4.2 29.8 48.2 20.2 1.8 114
Google Gemma 3 3.4 26.1 10.4 63.5 115
DeepSeek-VL 32.8 36.2 38.8 11.2 13.8 80
Qwen2-VL 35.3 42.9 39.0 1.3 16.8 77
Closed Source
GPT-4o 36.1 39.5 43.4 17.1 76
Gemini 2.5 Flash-Lite 41.2 51.4 21.4 5.7 21.5 70
Claude Sonnet 4 52.1 45.6 28.1 26.3 57
Results. Step-level translation is locally tractable: models often produce syntactically valid lines, but compilation failures reveal semantic underspecification. Claude achieves the highest valid rate (52.1%), while DeepSeek-VL and Qwen2-VL are the strongest among open-source models. Errors are dominated by syntax (brackets, references) and undefined stitches, underscoring the difficulty of maintaining state consistency even when local syntax is correct.
3.4.2 Project-Level Translation
In the project-level setting, the model is provided with the complete crochet instruction in natural language together with the corresponding product image, and must generate an entire CrochetPARADE program. This variant is globally self-contained but considerably more challenging than the step-level task: models must track stitch states over long horizons, resolve ambiguities in natural language, and produce code that is both syntactically valid and semantically aligned with the final design. Image grounding plays a crucial role in disambiguating constructs such as repeated motifs, symmetry, and termination conditions.
Evaluation.
We assess model outputs using two complementary metrics. The first is Compilation Success Rate (CSR), and the second is Partial Executable Rate (PER), which measures the average fraction of a program that compiles successfully before failure. While CSR captures all-or-nothing executability, PER provides a finer-grained view of structural alignment, offering credit to models that generate correct prefixes even if the full program does not compile. Error types are categorized using the same taxonomy as in the step-level evaluation.
Table 7: Project-level verification results. We report Compilation Success Rate (CSR) and Partial Executable Rate (PER). Best scores are bold, second-best are underlined.
Model Size CSR (%) PER (%)
Open Source Salesforce BLIP-2 Flan-T5 XL 3B 1.0 0.00
Google Gemma 3 4B 1.6 5.29
DeepSeek-VL 7B 8.1 37.49
Qwen2-VL 7B 21.0 30.28
Closed Source GPT-4o 4.0 2.76
Gemini 2.5 Flash-Lite 4.0 5.67
Claude Sonnet 4 5.0 8.16
Table 8: Project-level error type distribution across models (percentage). Abbreviations: Undef. (Undefined), Br. (Brackets), Lbl. (Labels Missing), Non-adj. (Non-adjacent), MRef (Multi-Reference), Other (Other/Runtime).
Model Undef. Br. Lbl. Non-adj. MRef Other
Open Source
Salesforce BLIP-2 Flan-T5 XL 2.0 37.4 58.6 2.0
Google Gemma 3 39.3 14.8 11.5 8.2 14.8 11.5
DeepSeek-VL 36.3 45.1 4.4 11.0 3.3
Qwen2-VL 25.0 12.5 7.8 51.6 3.1
Closed Source
GPT-4o 61.5 7.3 16.7 4.2 4.2 6.2
Gemini 2.5 Flash-Lite 51.1 18.1 9.6 8.5 4.3 8.6
Claude Sonnet 4 46.3 14.7 21.1 7.4 5.3 5.3
Results. Project-level translation remains highly challenging. Qwen2-VL achieves the best valid rate (21.0%) and strong PER, surpassing all closed-source systems. DeepSeek-VL demonstrates robust partial executability despite a lower valid rate. By contrast, GPT-4o, Gemini, and Claude achieve lower scores, highlighting that even state-of-the-art VLMs struggle with long-range structural consistency in executable synthesis. Error analysis (Table 8) reveals that closed-source models often fail due to label and reference inconsistencies, while open-source models more frequently exhibit undefined stitches and syntax errors.
4 Discussion
The results across CrochetBench highlight both the promise and current limitations of Visual language models in bridging perception, procedural reasoning, and executable synthesis. By structuring tasks in a progressive ladder, we expose clear gradients of difficulty: while contemporary models demonstrate competence in low-level recognition and mid-level comprehension, their performance declines substantially when asked to generate or formalize instructions into compilable domain-specific programs. This pattern underscores both methodological bottlenecks in multimodal grounding and fundamental challenges in symbolic reasoning over long-horizon structures.
In Task A, stitch recognition is feasible for both open- and closed-source models, though precision–recall tradeoffs differ. Claude Sonnet 4 prioritizes precision for a higher F1, while DeepSeek-VL favors recall via richer visual encoders. This reflects divergent inductive biases: commercial models better regularize spurious predictions, while open-source models overpredict to capture subtle textures. Crucially, this is a multi-label, fine-grained texture classification task with symbolic implications, distinguishing it from generic object recognition. Baselines like BLIP-2, designed for captioning, underperform due to insufficient symbolic grounding (blip; clip). Task B amplifies these challenges. Closed-source models, especially GPT-4o, outperform open models by a wide margin, underscoring the difficulty of aligning visual cues with procedural semantics. Distractor instructions are intentionally plausible, demanding reasoning over local–global coherence, not just token overlap (GVE; VQA; miech2019howto100mlearningtextvideoembedding). Qwen2-VL’s competitive performance suggests that scale and pretraining diversity help, but commercial systems benefit from stronger instruction tuning and better vision–language alignment.
Task C exposes the steep drop in open-ended procedural generation. Gemini 2.5 leads across BLEU, ROUGE, and ChrF (papineni2002bleu; lin2004rouge; popovic2015chrf), reflecting fluency in structured text. Yet absolute scores are low, and outputs often fail to conform to crochet-specific syntax. Open models hallucinate frequently, lacking exposure to domain-aligned distributions. This highlights a broader issue: fluency in general language modeling does not imply competence in structured, domain-specific generation (miech2019howto100mlearningtextvideoembedding; alayrac2022flamingovisuallanguagemodel).
Task D proves most demanding. Even top closed-source models rarely exceed 6% project-level compilation success, while Qwen2-VL unexpectedly achieves 21%, suggesting better symbolic generalization under execution constraints. Syntax errors (e.g., undefined stitches, malformed brackets) dominate in open models, while closed models produce syntactically valid but semantically inconsistent programs. These complementary failure modes hint at differing generalization paths: symbolic robustness versus lexical fluency. Notably, Task D also tests 3D-aware procedural fidelity—models must translate visual or textual cues into symbolic programs that unfold into spatially coherent structures, not just grammatically valid sequences.
Taken together, these results demonstrate that success in linguistic generation does not translate directly to executable synthesis. Performance decays sharply when evaluation moves from surface-level fidelity (BLEU, ROUGE) to structural validity (compilation), reinforcing the importance of execution-grounded metrics for procedural tasks. The relative strength of Qwen2-VL at the project level further suggests that progress may come not from scaling alone, but from architectural or training adjustments that better capture long-range dependencies and stateful operations. More broadly, CrochetBench exposes a critical frontier for multimodal reasoning: models must not only ground text in vision but also internalize procedural invariants that guarantee functional correctness.
However, our analysis must be contextualized within certain constraints. The benchmarks reflect standardized stitch sets and normalized instructions, which, while ensuring comparability, simplify the variability encountered in real-world crochet practice (e.g., designer-specific shorthand, unconventional repeats). Moreover, evaluation metrics in Task C rely on string overlap, which may underestimate semantically correct but lexically divergent outputs. Even the compilation-based metric in Task D, though stronger, cannot assess visual fidelity unless paired with image-render verification. Finally, the sample size for project-level DSL translation remains limited, and absolute validity rates are low, constraining fine-grained statistical comparisons.
5 Future Work
A central avenue for future research is advancing models that translate free-form natural language crochet instructions into the formal CrochetPARADE DSL. As a domain-specific programming language, CrochetPARADE not only enables executable verification of multimodal outputs but also positions crochet as a program synthesis problem, where compilers map symbolic grammars into machine-executable instructions, extending ideas from prior DSL work in domains such as knitting (knitscript), graphics (ellis2018learninginfergraphicsprograms), and robotics (mu2024robocodexmultimodalcodegeneration). This perspective connects naturally to CAD/CAM integration: industrial crochet and warp-knitting machines (e.g., COMEZ, Jakob Müller) already rely on pipelines from graphical design interfaces, to intermediate graph-based representations, to low-level machine code. CrochetPARADE could serve as a standardized intermediate layer in this workflow, bridging human-facing authoring tools with machine-facing execution systems (text2cad).
Several technical directions arise from this framing. First, incorporating explicit state-tracking mechanisms—whether through memory-augmented architectures (graves2016hybrid) or symbolic scaffolds (nye2021workscratchpadsintermediatecomputation)—could mitigate long-range inconsistencies in DSL translation. Second, multimodal pretraining enriched with procedural and topological domains (e.g., assembly instructions (cao2023instruction), instructional videos (miech2019howto100mlearningtextvideoembedding)) may narrow the gap between natural-language fluency and executable synthesis. Third, hybrid evaluation pipelines that combine compilation checks with visual render comparisons could more holistically assess structural and perceptual fidelity, building on metrics used in programmatic 3D generation (jain2022zero; paschalidou2021neuralpartslearningexpressive). Finally, CrochetBench offers a testbed for exploring neuro-symbolic integration, where neural perception is paired with symbolic reasoning to reconcile fine-grained visual cues with globally coherent program execution.
In sum, CrochetBench and CrochetPARADE highlight a pronounced frontier at the intersection of multimodal learning, program synthesis, and digital fabrication. Progress here may one day close the loop from human intent to automated textile production, advancing both structured multimodal reasoning and computational craft.
Ethics Statement
We acknowledge that the original crochet pattern PDFs are protected under copyright and therefore do not distribute raw files or full texts. Instead, we release only structured JSON annotations generated with GPT, reference URLs to the original sources, and our parsing and annotation scripts. The benchmark is provided strictly for non-commercial academic use. This approach enables reproducible research while respecting intellectual property and ensuring that our dataset serves as a tool for studying structured generation rather than redistributing creative works.
Reproducibility Statement
We have taken several steps to ensure the reproducibility of our results. All datasets, task templates, and evaluation procedures are documented in the main text and appendix. An anonymous repository containing the full source code, experiment scripts, and detailed reproduction instructions has been made publicly available at: https://github.com/Peiyu-Georgia-Li/crochetBench. This ensures that all reported results can be independently verified and extended by the research community.
Appendix A Related Work
A.1 Multi-modal Datasets Beyond Captioning
Most multimodal benchmarks have centered on descriptive pairing of images and natural language. Large-scale resources such as COCO [lin2015microsoftcococommonobjects] and Flickr30k [plummer2016flickr30kentitiescollectingregiontophrase] provide dense captions of everyday scenes, advancing vision–language representation learning. More recent datasets extend beyond captioning to procedural or instructional domains. Recipe1M+ aligns food images with ingredient lists and cooking steps [marin2019recipe1mdatasetlearningcrossmodal], while instructional video corpora such as YouCook2 [zhou2017automaticlearningproceduresweb] and HowTo100M [miech2019howto100mlearningtextvideoembedding] pair narrated demonstrations with visual segments. These resources emphasize semantic alignment but generally evaluate with retrieval- or similarity-based metrics.
Our benchmark departs from this paradigm by pairing images with executable procedures. Rather than asking models to generate a semantically similar description, we require them to synthesize a program (CrochetPARADE DSL) that can be rendered and structurally verified. This shift enables functional evaluation—akin to program synthesis—and reduces reliance on subjective similarity measures.
A.2 Bridging Visual Reasoning and Procedural Language
Procedural understanding benchmarks highlight the importance of sequential, state-dependent reasoning. Datasets such as Assembly101 [sener2022assembly101largescalemultiviewvideo] and IKEA-ASM [benshabat2023ikeaasmdatasetunderstanding] capture human assembly activities, modeling dependencies across actions, objects, and preconditions. Instructional video benchmarks further test long-horizon understanding and error detection [wei2025instructionbenchinstructionalvideounderstanding, lin2024videoguibenchmarkguiautomation]. Our task complements this line of work by grounding supervision not in temporally segmented actions but in artifact-centric procedures—finished crochet items paired with symbolic, stepwise instructions. This enables models to reason about topology, geometry, and sequential dependencies in a single unified representation.
A.3 Crafting and Domain-Specific Procedural Data
Closer to our domain, prior work has begun to explore fiber crafts. Seitz et al. introduced Digital Crochet, a visual, graph-based notation system for representing crochet patterns [digital_crochet]. In knitting, Kaspar et al. developed Neural Inverse Knitting, mapping images of knitted patterns to machine instructions [kaspar2019neuralinverseknittingimages]. These works underscore the feasibility of executable supervision in crafts, but remain limited in scale and scope. Our benchmark builds on this foundation by providing thousands of real, community-tested crochet patterns, along with a compiler and renderer for executable evaluation.
A.4 Relation to Code Benchmarks
The executable nature of CrochetPARADE connects it directly to program synthesis benchmarks. HumanEval [chen2021evaluatinglargelanguagemodels], MBPP [austin2021programsynthesislargelanguage], and APPS [hendrycks2021measuringcodingchallengecompetence] evaluate code generation by execution against unit tests. Similarly, Spider [yu2019spiderlargescalehumanlabeleddataset] and CoNaLa [yin2018mining] frame natural language to code translation tasks. In vision, Im2LaTeX-100K [deng2017imagetomarkupgenerationcoarsetofineattention] and pix2code [beltramelli2017pix2codegeneratingcodegraphical] evaluate image-to-program translation with render fidelity as the metric. Our tasks—ranging from text-to-DSL translation to image-to-stitch recognition—extend this paradigm to the crafting domain, introducing structured 2D/3D topology as part of the evaluation. This grounds multimodal learning in a setting where success requires both semantic alignment and structural correctness.
A.5 Domain-Specific Languages for Executable Evaluation
A key enabler of our benchmark is the use of a domain-specific language. Prior DSL-based benchmarks, such as Im2LaTeX and the CROHME competition on handwritten math recognition [crohme], demonstrate how symbolic formalisms enable reproducible rendering and structural evaluation. CrochetPARADE adopts this principle for fiber crafts: each pattern compiles into an abstract syntax tree that can be rendered and tested, supporting metrics such as ExecPass@K, structural unit tests, and image–render similarity. This functional perspective moves beyond surface-level similarity to test whether a model’s output actually works, in contrast to chemistry and recipe domain, which require quite amount of time to validate it [huang2024application, chefFusion, huang2025chemhgnn].
Appendix B Additional Dataset Statistics
B.1 Instruction Complexity by Skill Level
Skill Level Avg. Length Median Length Avg. Abbr. Count
Beginner 1,674 1,365 9.2 465
Easy 2,761 2,182 10.8 3,569
Intermediate 4,221 3,387 10.7 1,967
Experienced 7,689 6,729 9.8 80
Table 9: Instruction complexity by skill level. Length is measured in characters.
Example Dataset Entry
Table 10: Representative pattern entry from CrochetBench.
Field Value
Pattern Name SKULL TRICK OR TREAT BAG (TO CROCHET)
Skill Level Intermediate
Project Type Bags or Purses
Measurements 15 cm diameter $\times$ 15 cm high (excluding handle)
Gauge 13 sc and 14 rows = 10 cm
Materials Lily® Sugar’n Cream (White, Black), 5 mm hook, cardboard
Image https://www.yarnspirations.com/cdn/shop/products/SCC0303-005314M.jpg
Source input_file/Bags+Purses/SCC0303-005314M.pdf
Instructions (truncated for brevity)
B.2 Skill Level Distribution
B.2.1 Overall Distribution
Table 11 summarizes the overall distribution of skill levels across the CrochetBench dataset. The majority of patterns are labeled as easy (58.7%), followed by intermediate (32.3%). Only a small fraction are classified as beginner (7.6%) or experienced (1.3%). [^3] Three additional rare labels were observed: easy to intermediate (1 pattern), beginners (1 pattern), and beginner/easy (1 pattern). Together they account for $0.1\%$ of the dataset.
Table 11: Overall skill level distribution. Percentages are relative to all patterns with annotated skill levels.
Skill Level Count Percentage
Easy 3569 58.66%
Intermediate 1967 32.33%
Beginner 465 7.64%
Experienced 80 1.31%
Total 6084 100%
One pattern (0.02%) is missing an annotated skill level.
B.2.2 Distribution by Project Type
We further break down skill levels by the top 10 most common project types. Results are shown in Table 12. In most categories, easy patterns dominate, typically ranging between 53–70%. Intermediate is the second most common, while beginner and experienced remain consistently low across categories.
Table 12: Skill level distribution by top 10 project types. Percentages are within each project category.
Project Type Easy Intermediate Beginner Experienced
Afghans & Blankets 56.1% 35.3% 7.0% 1.5%
Hats 61.3% 27.8% 10.1% 0.7%
Sweaters & Cardigans 56.6% 35.9% 5.0% 2.5%
Shawls & Wraps 52.7% 41.8% 4.2% 1.2%
Scarves 63.2% 20.7% 16.1%
Pillows & Poufs 70.0% 22.9% 6.5% 0.7%
Amigurumi & Toys 64.0% 33.2% 2.1% 0.7%
Bags & Purses 53.8% 39.0% 6.8% 0.4%
Décor 58.4% 33.3% 6.5% 1.7%
Dishcloths 62.6% 27.5% 9.9%
Overall, the predominance of easy patterns reflects the accessibility of crochet as a craft and aligns with the goal of many project types to cater to a wide audience. The relative scarcity of experienced-level patterns suggests that most published resources emphasize broad usability rather than advanced expertise.
B.3 Pattern Complexity Analysis
B.3.1 Instruction Length Statistics
We first analyze the distribution of instruction lengths, measured in raw character counts. As shown in Table 13, the average instruction length is over 3,200 characters, while the median is substantially lower at 2,453 characters, reflecting a long-tailed distribution. The most complex patterns extend beyond 30,000 characters, while some very short patterns are as small as 20 characters.
Table 13: Instruction length statistics (in characters).
Statistic Value
Average 3216.0
Median 2453.0
Min 20
Max 30634
25th percentile 1511.8
75th percentile 4136.2
90th percentile 6403.9
Out of 6,085 total patterns, 6,084 (99.98%) contain full instructions.
B.3.2 Abbreviation Statistics
Abbreviations, such as sc, dc, and hdc, are a distinctive element of crochet instructions. Table 14 reports abbreviation counts across all patterns. Most patterns contain about 10 abbreviations, with values ranging from 1 to 31.
Table 14: Abbreviation count statistics.
Statistic Value
Average 10.6
Median 10.0
Min 1
Max 31
B.3.3 Complexity by Skill Level
Instruction length correlates with the designated skill level. As shown in Table 15, beginner-level patterns average under 2,000 characters, while intermediate patterns extend to over 4,200. Experienced patterns are the longest, averaging 7,689 characters. Rare categories such as easy to intermediate skew extremely long due to outliers.
Table 15: Instruction length and abbreviation counts by skill level.
Skill Level Avg. Length Median Length Avg. Abbr. Count
Easy to intermediate 13812.0 13812.0 21.0 1
Experienced 7689.4 6729.0 9.8 80
Intermediate 4221.3 3387.0 10.7 1967
Easy 2760.7 2182.0 10.8 3569
Beginner 1673.9 1365.0 9.2 465
Beginners 1633.0 1633.0 11.0 1
Beginner/Easy 1063.0 1063.0 1
B.3.4 Most and Least Complex Project Types
Finally, we identify the most complex and simplest project types by average instruction length. Tables 16 and 17 list the top 10 categories. Garments such as dresses, vests, pants, and tunics are the most demanding, with average instructions exceeding 5,800 characters. By contrast, smaller accessories such as cowls, washcloths, scarves, and headbands are substantially shorter, typically under 2,000 characters.
Table 16: Top 10 most complex project types (by average instruction length).
Project Type Avg. Length Median Count
Dresses 6484.9 5799.0 34
Vests 6032.0 5193.5 64
Pants 5866.7 5409.0 11
Tunics 5850.4 5832.0 29
Sets 5625.5 4847.0 111
Sweaters & Cardigans 5429.2 5113.0 357
Amigurumi & Toys 5322.4 4505.0 286
Jackets 5311.9 4831.0 31
Onesies & Rompers 5263.4 5181.0 5
Aprons 4467.8 4494.0 11
Table 17: Top 10 simplest project types (by average instruction length).
Project Type Avg. Length Median Count
Cowls 1288.3 956.5 154
Washcloths & Mitts 1502.5 1420.0 28
Scarves 1567.3 1221.0 304
Headbands 1617.5 1475.5 38
Dishcloths 1688.4 1571.0 222
Coasters 1750.3 1625.0 26
Booties 1921.9 1938.5 24
Jewelry 1960.3 1549.0 55
Super Scarves 2007.6 1213.0 13
Tech Accessories 2011.1 2099.0 13
Taken together, these results highlight strong alignment between project type, designated skill level, and instruction length. Garment-oriented projects require substantially longer and more complex instructions, while accessories and small decorative items remain simple and concise.
Appendix C Prompts
C.1 Task A: Stitch Recognition Prompt
This task evaluates a model’s ability to identify stitches present in a crochet product image.
Stitch Recognition Prompt (Rendered Example) SYSTEM PROMPT You are a crochet stitch expert. Given an image of a crochet product, identify all stitches that appear. Requirements: - Use only standard U.S. crochet abbreviations (e.g., sc, hdc, dc, tr, ch, sl st, pop, etc.). - Output must be a comma-separated list of abbreviations. - Do not include explanations, extra text, or formatting beyond the list. USER PROMPT Look at this crochet product image and list the stitches used. [Image]
C.2 Task B: Instruction Selection Prompt
This task evaluates a model’s ability to choose the correct instructions from multiple-choice options.
[…]
C.3 Task C: Instruction Generation Prompt
This task evaluates a model’s ability to generate complete crochet instructions from an image.
Instruction Generation Prompt (Rendered Example) SYSTEM PROMPT You are a professional crochet pattern writer. Examine the image of the finished crochet product carefully. Write a complete set of crochet instructions in the standard style used in published patterns. Requirements: - Use standard abbreviations: sc (single crochet), hdc (half double crochet), dc (double crochet), tr (treble), ch (chain), sl st (slip stitch), rep (repeat). - Organize the instructions row by row or round by round (e.g., ”Rnd 1: …”, ”Row 2: …”). - If color changes are visible in the image, include them in the pattern. - Keep the instructions concise and precise, as if for experienced crocheters. - Output only the crochet pattern. Do not add any explanations, commentary, or extra text. USER PROMPT Generate step-by-step crochet instructions for this image. [Image]
C.4 Task D (Step-level): NL $\rightarrow$ DSL Translation Prompt
This task evaluates whether a model can translate a single natural language instruction into exactly one line of compilable CrochetPARADE DSL code.
Step-level NL $\rightarrow$ DSL Translation Prompt (Rendered Example) SYSTEM PROMPT You are a crochet compiler. Translate the next instruction NL into one line of CrochetPARADE DSL. Use consistent naming and syntax. Important rules for translations: 1. Make sure your output ONLY contains the DSL code, nothing else. 2. Use the previous examples to understand the pattern of translation. 3. Be consistent in naming conventions with the examples. 4. Your output should be exactly one line of DSL code. USER PROMPT Now translate the NL into DSL: NL: DSL:
C.5 Task D (Project-Level): NL $\rightarrow$ DSL Translation Prompt
This task evaluates whether a model can convert natural language crochet instructions (with optional images) into compilable CrochetPARADE DSL code.
NL $\rightarrow$ DSL Translation Prompt (Rendered Example) SYSTEM PROMPT You are a professional crochet pattern writer. Convert instructions + images into compilable CrochetPARADE DSL code. Output only the DSL code. No explanations, commentary, or extra text. Example 1: ”image path”: https://www.yarnspirations.com/cdn/shop/files/BRC0116-035467M.jpg, INSTRUCTIONS Note: Join with sl st to first sc at end of each rnd. Ch 2. **Rnd 1:** 6 sc in 2nd ch from hook. Join. (6 sc) **Rnd 2:** Ch 1. 2 sc in each sc around. Join. (12 sc) **Rnd 3:** Ch 1. (2 sc in next sc, 1 sc in next sc) repeat around. End with 1 sc. Join. (18 sc) **Rnd 4:** Ch 1. (2 sc in next sc, 1 sc in each of next 2 sc) repeat. End with 1 sc in last 2 sc. Join. (24 sc) **Rnd 5:** Ch 1. Sc in each sc around. Join. (24 sc) **Rnd 6:** Ch 1. (2 sc in next sc, 1 sc in each of next 3 sc) repeat. End with 1 sc in last 3 sc. Join. (30 sc) **Rnds 7–8:** Repeat Rnd 5 (sc in each sc). Join. (30 sc each round) **Rnd 9:** Ch 1. **Working in back loops only**: (2 sc in next sc, 1 sc in each of next 2 sc) repeat. End with 1 sc in last 2 sc. Join. (40 sc) **Rnd 10:** Ch 1. Sc in each sc around (both loops). Join. (40 sc) **Rnd 11:** Ch 1. (2 sc in next sc, 1 sc in each of next 3 sc) repeat. End with 1 sc in last 3 sc. Join. (50 sc) **Finish:** Fasten off. DSL ¶ch.B ¶sc@B.A,5sc@B,ss@A ¶ch.A,sk,6sc2inc,ss@A ¶ch.A,sk,[sc2inc,sc]*6,ss@A ¶ch.A,sk,[sc2inc,2sc]*6,ss@A ¶ch.A,sk,24sc,ss@A ¶ch.A,sk,[sc2inc,3sc]*6,ss@A ¶[ch.A,sk,30sc,ss@A ¶]*2 ¶ch.A,sk,[scbl,scbl@[@],2scbl]*10,ss@A ¶ch.A,sk,40sc,ss@A ¶ch.A,sk,[sc2inc,3sc]*10,ss@A USER PROMPT Now generate DSL code for the following: [Image] [Instructions] Rnd 1: Ch 2, 6 sc in ring Rnd 2: 2 sc in each (12) Rnd 3: [Sc, sc, inc] around (16) Rnd 4: [Tr, sc] repeat around [DSL]
Appendix D CrochetPARADE: Pattern Renderer, Analyzer, and Debugger
CrochetPARADE (short for Crochet Pattern Renderer, Analyzer, and Debugger) is an interactive platform that enables users to author, visualize, test, and export crochet patterns in both 2D and 3D [crochetparade]. By combining a custom pattern grammar with simulation and rendering tools, CrochetPARADE addresses common issues of ambiguity, correctness, and interpretability in textual crochet instructions. [^4] https://www.crochetparade.org/
Core Capabilities.
  • Interactive authoring and rendering. Users write pattern instructions in the CrochetPARADE grammar and then invoke a “calculate” operation to convert those instructions into a virtual model. The system supports both 2D and 3D views, along with interactive controls such as zoom, rotation, and stitch highlighting.
  • Validation and debugging. CrochetPARADE parses the input, checks for syntactic and consistency errors (e.g., mismatched stitch counts, impossible attachments), and flags over- or under-stretched stitches.
  • Export and interoperability. From a rendered pattern, users can export:
    • A standard crochet chart (SVG) with conventional stitch symbols and labeled stitch connections.
    • A 3D model (GLTF format) for integration into external tools such as Blender.
    • The underlying pattern instructions text (in the CrochetPARADE grammar), ensuring reproducibility and sharing.
Design Ideals and Rationale.
CrochetPARADE is built to meet several design goals: (i) unambiguous precision, where the grammar is far more strict than free-form natural language, reducing interpretive errors; (ii) local computation, since all parsing, simulation, and rendering occur client-side in the browser with no user instructions sent to a central server; and (iii) open source extensibility, as the platform is released under GPLv3, with the grammar manual provided under a Creative Commons BY-NC-SA license.
Role in Our Work.
Within the context of CrochetBench, CrochetPARADE provides a rigorous target representation: model predictions can be compiled into CrochetPARADE instructions, validated for syntactic and structural correctness, and then visualized or executed. This enables evaluation beyond surface-level metrics (e.g., BLEU, ROUGE) toward executor correctness—whether a generated pattern is valid, renderable, and stitch-balanced.
Figure 3: Example of the CrochetBench translation pipeline. (Left) Natural language crochet instructions from the dataset. (Second) Automatically translated into CrochetPARADE DSL, a formal stitch grammar. (Third) Mesh rendering generated from the DSL. (Right) Target crocheted item image provided in the dataset. This pipeline enables direct text-to-image consistency checks, automated validation, and future training of NL $\rightarrow$ DSL models, analogous to text-to-code generation.
Appendix E DSL Error Taxonomy
To better understand failure cases in Task D, we extend the validator’s error analysis with detailed subcategories and examples. Errors fall into four major groups:
1. Syntax Structure Errors
Unbalanced Brackets.
Missing opening/closing parentheses or brackets.
Examples Unbalanced brackets: (sc,hc5,sltr)infl)
Multiple References Without Parentheses.
Improper formatting of references.
Example Multiple references defined without parenthesis: (21ch),turn
sk,(20sc)
(2ndrow):Ch1.(1scbl)ineachchtoendofrow.Turn
2. Stitch Definition Errors
Stitch Not Found.
Undefined stitch types not in the dictionary.
Examples ch1, ch3, scfp, hdc_bar
3. Labeling and Reference Errors
Label Not Found.
Reference to a non-existent label.
Example Label not found: C
Reusing Labels Incorrectly.
Same label used for non-adjacent stitches.
Example Cannot use same label over non-adjacent stitches. Consider using different labels.
4. Structural and Formatting Issues
Turning Errors.
Misplaced turning commands.
Example Turning can happen only at the end of a row.
Variable Naming Issues.
Conflicts between variable names and stitch names.
Example Error: variable name matches stitch name. For example, \$ch=0\$ cannot be used since ’ch’ is a stitch name.
Runtime Errors.
Low-level parsing failures from the JavaScript compiler.
Examples Cannot read properties of null (reading ’0’) Cannot use ’in’ operator to search for ’attach_id’ in NaN

### v2

@@
 Abstract
-We present CrochetBench, a benchmark for evaluating the ability of multimodal large language models to perform fine-grained, low-level procedural reasoning in the domain of crochet.
-Unlike prior benchmarks that focus on high-level description or visual question answering, CrochetBench shifts the emphasis from describing to doing: models are required to recognize stitches, select structurally appropriate instructions, and generate compilable crochet procedures.
+While multimodal large language models can describe visual content, their ability to generate executable procedures remains underexplored.
+CrochetBench presented in this paper evaluates this shift from describing to doing through fine-grained procedural reasoning in crochet: models must recognize stitches, select structurally appropriate instructions, and generate compilable procedures.
 We adopt the CrochetPARADE DSL as our intermediate representation, enabling structural validation and functional evaluation via execution.
 The benchmark covers tasks including stitch classification, instruction grounding, and both natural language and image-to-DSL translation.
-Across all tasks, performance sharply declines as the evaluation shifts from surface-level similarity to executable correctness, exposing limitations in long-range symbolic reasoning and 3D-aware procedural synthesis.
-CrochetBench offers a new lens for assessing procedural competence in multimodal models and highlights the gap between surface-level understanding and executable precision in real-world creative domains.
-Code is available at https://github.com/Peiyu-Georgia-Li/crochetBench.
+Across all tasks, performance sharply decreases as the evaluation shifts from surface-level similarity to executable correctness, revealing limitations in long-range symbolic reasoning and 3D-aware procedural synthesis.
+Our proposed CrochetBench offers a new lens for assessing procedural competence in multimodal models and highlights the gap between surface-level understanding and executable precision in real-world creative domains.
+Code is available at https://anonymous.4open.science/r/crochet-82E6/README.md.
+CrochetBench: Can Vision-Language Models Move from Describing to Doing in Crochet Domain?
+Peiyu Li*, Xiaobao Huang*, Ting Hua, Nitesh V. Chawla University of Notre Dame Notre Dame, IN 46556 USA {pli9, xhuang2, thua, nchawla}@nd.edu
+**footnotetext: These authors contributed equally.
+Figure 1: End-to-end data construction and benchmarking workflow for CrochetBench.
+The left panel illustrates the data collection pipeline: we download PDF files and image links from crochet pattern websites, and parse them using GPT-4o-mini to produce structured JSON files.
+From each JSON record, we derive four supervised datasets: (A) stitch-level labels, (B) multiple-choice instruction selection items, (C) natural-language instruction generation pairs, and (D) expert-annotated DSL programs for procedural synthesis.
+The right panel summarizes the four benchmarking tasks: Task A evaluates multi-label visual stitch recognition; Task B measures vision-to-instruction grounding via MCQ selection; Task C assesses vision-conditioned natural-language instruction generation; and Task D tests stateful procedural reasoning via NL-to-DSL (Natural Language to Domain-Specific Language) translation with execution-based metrics.
 1 Introduction
-Procedural crafts such as crochet present a distinctive frontier for multimodal learning.
-Unlike traditional captioning or recipe datasets (chefFusion; Hu_2023_ICCV; mohbat2024llava), crochet patterns intertwine three interdependent modalities: (i) structured symbolic language, where stitch abbreviations and counts define a precise grammar of construction; (ii) long-form natural language, which provides contextual guidance such as materials and sizing; and (iii) visual evidence, including photographs of completed objects and motif diagrams.
-Success requires not just alignment across modalities but step-wise reasoning that preserves procedural fidelity, making the challenge closer to program synthesis than generic description.
-Crochet also offers a unique testbed for 3D-aware reasoning.
-Each stitch encodes both local geometry and global connectivity, forming a topological structure that must be preserved across steps.
-Generating or interpreting patterns thus demands reasoning over how sequential operations accumulate into volumetric form.
-In effect, crochet couples symbolic instruction following with embodied spatial reasoning, cultivating abilities essential for domains where language must ground into physical tasks.
-Despite the rapid growth of multimodal benchmarks (fu2024mmecomprehensiveevaluationbenchmark; li2023seed; zhang-etal-2025-ngqa; yue2024mmmu), existing datasets have largely focused on description or grounding.
-COCO (lin2015microsoftcococommonobjects) catalyzed captioning research, TextCaps (sidorov2020textcapsdatasetimagecaptioning) extended it to text-in-the-wild, and Recipe1M (marin2019recipe1mdatasetlearningcrossmodal) explored cross-modal cooking instructions.
-While recipes also involve multi-step procedures, validating correctness typically requires real-world execution, making large-scale evaluation slow and resource-intensive.
-Crochet, by contrast, provides a symbolic domain where outputs can be automatically verified through DSL compilation, enabling scalable and efficient study of step-wise reasoning.
-For a more detailed survey, see Appendix A.
-Yet these benchmarks stop short of testing whether models can follow symbolic grammars, respect numerical and spatial constraints, and produce outputs that are executable.
-Current systems can describe, but not reliably do.
-CrochetBench fills this gap by centering evaluation on instructional fidelity: can models not only recognize and generate, but also output step-wise, compilable instructions that respect symbolic, numerical, and topological structure?
-Each example in CrochetBench is a multimodal package—structured JSON metadata (stitch inventories and abbreviations), full-text procedures with rows/rounds and conditionals, and paired images of finished objects and motifs.
-Crucially, CrochetBench is paired with CrochetPARADE (crochetparade), a domain-specific language (DSL) enabling executable evaluation, where natural language instructions are translated into compilable code enforcing geometric and topological coherence.
-Our contributions are fourfold: (1) CrochetBench, the first executable benchmark for procedural crafts, unifying symbolic, textual, and visual modalities with evaluation protocols emphasizing procedural fidelity and 3D-aware reasoning; (2) a comprehensive task suite spanning recognition, comprehension, generation, and DSL translation; (3) integration of CrochetPARADE into an executable pipeline, enabling scalable, automated verification of outputs—unlike domains such as cooking, which require real-world execution—thereby shifting evaluation from surface similarity to procedural fidelity; and (4) baseline analyses of state-of-the-art VLMs/MLLMs, revealing systematic weaknesses including hallucinations, captioning bias, and structural artifacts.
-Taken together, CrochetBench opens a new direction for multimodal research: moving beyond describing what we see, toward generating executable procedures that respect symbolic grammar, numerical accuracy, and topological coherence—paving the way for models that can reason in structured 3D spaces.
+Multimodal large language models can describe what they see, but can they explain how to make it?
+Given an image of a piece of furniture or a crocheted blanket, current models can fluently describe its appearance, yet cannot generate the step-by-step instructions needed to assemble or recreate it.
+Closing this gap, from perception to executable procedure, is critical for applications in robotics, CAD design, and procedural content generation.
+However, evaluating procedural competence is difficult: most real-world procedures require physical execution to verify correctness.
+One domain offers a rare exception: crochet instruction synthesis.
+These crochet instructions follow a formal symbolic grammar that can be compiled and validated automatically, enabling scalable evaluation of procedural competence without requiring costly physical execution.
+Existing multimodal benchmarks largely focus on descriptive tasks.
+Image captioning benchmarks such as COCO (Lin et al., 2014) and TextCaps (Sidorov et al., 2020) evaluate whether models can generate natural language descriptions of visual scenes.
+Text-to-code benchmarks such as HumanEval (Chen, 2021a) evaluate program synthesis but lack visual grounding, while image-to-code benchmarks such as pix2code (Beltramelli, 2017) test layout generation rather than stateful procedural reasoning.
+Even procedural datasets such as Recipe1M+ (Marin et al., 2018) primarily assess cross-modal alignment rather than executable correctness, since validating a recipe requires cooking it, a slow and resource-intensive process.
+No existing benchmark combines visual understanding, multi-step procedural decomposition, and automated executable validation.
+To address this gap, we turn to crochet: an ideal testbed for visual procedural reasoning.
+Each crochet pattern presents a vision-to-procedure challenge.
+Given an image of a finished artifact, models must infer the sequential construction process: which stitches were used, in what order, and with what counts.
+This requires (1) fine-grained visual recognition of stitch types from texture and topology, (2) procedural decomposition into ordered operations with symbolic grounding, (3) stateful reasoning to track counts across steps, and (4) constraint satisfaction to ensure topological consistency.
+What makes crochet unique is that, crochet instructions can be formalized into domain-specific languages and validated through compilation, which enables the automated executable evaluation that other procedural domains lack.
+In this paper, we introduce CrochetBench, the first benchmark for executable visual procedural reasoning in creative domains.
+Our CrochetBench comprises 6,085 crochet patterns spanning 55 categories with over 98% image coverage, ranging from beginner to expert complexity.
+We design four tasks of increasing difficulty: Task A (Stitch Recognition) evaluates multi-label visual classification; Task B (Instruction Selection) evaluates vision-to-language grounding via multiple-choice matching; Task C (Instruction Generation) evaluates free-form instruction synthesis from images; and Task D (DSL Translation) evaluates translation into executable CrochetPARADE programs (Tassev, 2025).
+Together, these tasks shift evaluation from surface-level fluency to procedural correctness, whether outputs compile and preserve structural constraints.
+We evaluate nine state-of-the-art vision-language models on CrochetBench, including GPT-4o (Hurst et al., 2024), Gemini 2.5 Flash-Lite (Comanici et al., 2025), Claude Sonnet 4 (Anthropic, 2025), and open-source alternatives spanning 3B to 72B parameters.
+Our evaluation reveals that current models exhibit a consistent gap between perception and procedural synthesis.
+Models achieve moderate performance on stitch recognition and instruction selection, but struggle to generate multi-step instructions or executable DSL programs.
+A case study (Figure 2) illustrates this pattern: even when models produce fluent instructions that correctly identify local details such as colors and motifs, the resulting procedures misconstruct global geometry and fail under execution.
+Also, neither model scaling nor supervised finetuning closes this gap.
+Larger models often achieve lower valid pattern rates during execution due to a tendency to invent novel stitch types not defined in the DSL, while finetuning improves text-based metrics without improving executable correctness.
+Our contributions are threefold: (1) We release CrochetBench, the first executable benchmark for procedural textile crafts, comprising 6,085 patterns integrated with the CrochetPARADE DSL for automated validation beyond text similarity.
+(2) We design four tasks of increasing difficulty—stitch recognition, instruction selection, instruction generation, and DSL translation—that enable fine-grained diagnosis of procedural competence.
+(3) We evaluate nine state-of-the-art vision–language models on CrochetBench and reveal a consistent gap between recognition and executable synthesis that persists across model scales and finetuning.
 2 Dataset Description
-Figure 1: Distribution of the top-10 most common project types in CrochetBench.
-CrochetBench is a large-scale, structured benchmark consisting of 6,085 crochet patterns spanning 55 distinct project categories.
-It was constructed by collecting publicly available patterns from the Yarnspirations website [^1] https://www.yarnspirations.com/collections/patterns, a widely used repository for fiber arts.
-The raw patterns were originally formatted as PDF documents, which were parsed and normalized using a GPT-4o-mini-based conversion pipeline.
-This process extracted and standardized key fields including pattern metadata, materials, measurements, gauge, abbreviations, and full step-by-step instructions.
-Each entry was converted into a machine-readable JSON object with a consistent schema, and 98.77% of patterns include an associated product image, enabling multimodal supervision.
-Figure 2: Skill level distribution across the CrochetBench dataset. Note that the “Experienced” slice (1.3%) is annotated externally due to its small size.
-The dataset supports diverse real-world crochet practices, with project types ranging from simple accessories to complex garments.
+CrochetBench is a large-scale, structured benchmark comprising 6,085 crochet patterns across 55 distinct project categories.
+As shown in the left panel of Figure 1, the dataset is constructed from publicly available patterns on the Yarnspirations website***https://www.yarnspirations.com/collections/patterns, a widely used repository in the fiber-arts community.
+The raw patterns (originally distributed as PDF documents) were parsed and normalized through a GPT-4o-mini–based conversion pipeline that extracted and standardized key fields such as metadata, materials, measurements, gauge, abbreviations, and full step-by-step instructions.
+Each pattern was then transformed into a JSON object following a consistent schema.
+Each pattern pairs a finished-product image with full natural-language instructions, enabling both perception- and generation-based evaluation.
+CrochetBench exhibits substantial diversity in both visual appearance and procedural complexity.
+Patterns vary widely in instruction length, stitch vocabulary, and structural depth, supporting evaluation across different levels of difficulty and long-horizon reasoning.
+Detailed statistics, skill-level distributions, and a JSON example are provided in the Appendix B.
+3 Tasks
+A central goal of CrochetBench is to evaluate whether multimodal LLMs can move beyond surface-level visual description and produce procedurally correct crochet instructions.
+As shown in Table 1, CrochetBench is organized as a progression of four tasks that isolate the core cognitive abilities required for real-world crochet reasoning.
+Tasks A and B focus on perception and comprehension, representing the minimum prerequisites for procedural understanding.
+Stitch recognition and instruction selection evaluate whether models can identify stitch types from images and correctly associate visual patterns with corresponding natural-language instructions.
+However, accomplishing these two tasks does not guarantee the ability to synthesize a valid crochet procedure.
+Tasks C and D therefore target procedural generation and formalization, requiring models to produce coherent, stepwise natural-language instructions or executable CrochetPARADE programs.
+These tasks demand the integration of visual grounding, temporal consistency, symbolic manipulation, and domain-specific constraints.
+The following subsections describe each task in detail.
+ID Ability Tested Task Evaluation Metrics Test Size Eval Set
+A Recognition Stitch Recognition F1, Precision, Recall 6,009 CrochetBench-A
+B Comprehension Instruction Selection Accuracy 6,003 CrochetBench-B
+C Generation Instruction Generation BLEU, ROUGE, ChrF 6,009 CrochetBench-C
+D Formalization Instr.-to-DSL (Step) Valid Pattern Rate 119 CrochetBench-Dstep
+Instr.-to-DSL (Project) Valid Pattern Rate, Dino Similarity 100 CrochetBench-Dproj
+Table 1: CrochetBench task overview and evaluation metrics.
+3.1 Task A: Stitch Recognition
+Task A evaluates a model’s ability to identify crochet stitch types from an image of a finished product.
+We construct CrochetBench-A, a subset of 6,009 examples from the full benchmark, where each product image is paired with ground-truth stitch annotations.
+These labels are derived from the official pattern instructions and normalized into a standardized set of stitch abbreviations (e.g., sc, hdc, dc) to ensure consistency across patterns.
+Unlike standard image classification, this is a multi-label prediction problem: multiple stitches may co-occur within the same image, often with subtle visual differences in texture and geometry.
+This task therefore probes fine-grained visual grounding of structured crochet semantics.
+Evaluation.
+For each example, we compute overlap between the predicted and reference stitch sets.
+True Positives (TP) are stitches correctly predicted; False Positives (FP) are stitches predicted but not in the reference; and False Negatives (FN) are stitches in the reference but missed by the model.
+From these counts, we compute precision (fraction of correct predictions among all predictions), recall (fraction of ground-truth stitches recovered), and F1 score (harmonic mean) Powers (2020).
+Metrics are averaged across examples to provide overall performance.
+This formulation rewards models that recover all present stitches while avoiding spurious predictions.
+Accurate stitch recognition is foundational for the benchmark, as later tasks (e.g., instruction selection and instruction generation) depend on robust detection of stitch primitives.
+3.2 Task B: Instruction Selection
+Task B evaluates whether a model can correctly associate an image of a finished crochet product image with its corresponding natural-language instruction.
+We construct CrochetBench-B, a subset of 6,003 examples, where each instance contains one ground-truth instruction and three distractor instructions sampled from the same project category (e.g., hats, rugs).
+Because distractors originate from the same category, they share similar visual and lexical structure, thereby increasing task difficulty and preventing solutions based on superficial lexical overlap.
+The answer distribution across options is approximately uniform (A: 24.9%, B: 25.7%, C: 23.7%, D: 25.7%), ensuring no positional bias.
+Evaluation.
+To support scalable and reproducible benchmarking, we formulate the task as a four-way multiple-choice question (MCQ).
+The model must select one option (A–D), with exactly one correct answer.
+Predictions are extracted using a deterministic regex-based parser that identifies explicit letter-based responses (e.g., “A”, “Option B”, “The answer is D”).
+Accuracy is used as the evaluation metric.
+This task provides a controlled measure of visual grounding and semantic alignment between images and procedural text, without requiring free-form generation.
+By forcing discrimination among near-neighbor instructions, Task B probes whether models can leverage fine-grained visual cues and domain-specific stitch semantics, which are essential precursors to reliable procedural instruction generation.
+3.3 Task C: Instruction Generation
+Task C evaluates a model’s ability to generate natural-language crochet instructions from an image of a finished item.
+We construct CrochetBench-C, a subset of 6,009 examples in which each image is paired with the corresponding ground-truth textual instruction.
+In contrast to captioning or stylistic description, this task requires generating a sequence of domain-specific commands (e.g., “Rnd 1: ch 4, 6 sc in ring”), each of which encodes precise stitch operations, counts, and ordering.
+Because real crochet patterns may include tens of steps, hierarchical structure (rounds, rows, substeps), and long-range dependencies, this task assesses whether models can infer the underlying procedural logic implied by the final visual product.
+The generated text must maintain consistent stitch semantics, preserve temporal ordering, and follow established formatting conventions used by human crafters.
+Evaluation.
+We evaluate generation quality using BLEU, ROUGE-L, and ChrF (Papineni et al., 2002; Lin, 2004; Popović, 2015), which together capture complementary aspects of textual fidelity in procedural instructions.
+BLEU measures overlap of word-level n-grams and thus reflects local lexical accuracy in stitch tokens and command sequences.
+ROUGE-L evaluates the longest common subsequence between the generated and reference patterns, capturing larger-scale ordering and structural alignment across multi-step procedures.
+ChrF operates on character-level n-grams, which makes it effective for crochet patterns where stitch abbreviations (e.g., sc, sc2tog) often differ by only a few characters.
+Word-based metrics treat such tokens as entirely distinct, whereas character-level comparisons can capture partial matches and small but semantically important variations.
+However, textual overlap metrics alone cannot reveal whether the generated instructions form a coherent or executable procedure.
+A model may generate instructions that appear fluent and pattern-like while still violating fundamental structural constraints, including inconsistent stitch counts, infeasible transitions, or unbalanced repeat constructions.
+To directly assess structural correctness and program-level understanding, we introduce Task D, which requires models to formalize correct natural-language instructions into a machine-checkable DSL representation.
+3.4 Task D: Instruction-to-DSL Translation
+Tasks A–C evaluate perception, retrieval, and natural-language generation, but they do not test whether a model can represent crochet procedures in a structured, machine-interpretable form.
+Crochet patterns are inherently programmatic: they contain loops, repeats, and counting logic that natural language expresses only implicitly, and that text-based metrics cannot reliably validate.
+Task D isolates this structural dimension by requiring models to translate correct natural-language instructions into an executable DSL, thereby revealing whether models grasp the underlying program-like structure of crochet.
+This capability is essential for true procedural reasoning, and we instantiate it using the CrochetPARADE DSL.
+We construct two variants of Task D: CrochetBench-Dstep (119 items) for step-level formalization and CrochetBench-Dproj (100 items) for project-level program synthesis.
+Step-Level Translation
+The step-level task evaluates incremental translation from natural-language crochet instructions to DSL, where each instruction updates a shared stitch state.
+In this setting, the model is given a prefix of correct NL–DSL pairs representing the pattern translated so far and must generate the corresponding DSL line for the next instruction.
+Because crochet patterns are stateful, correct translation requires maintaining consistency with previous steps, including stitch counts, repeat structures, turning logic, and round-to-round topology.
+This formulation tests whether models can map local textual cues into the structured symbolic operations of CrochetPARADE while preserving long-range procedural context.
+To capture variation in pattern progression, CrochetBench-Dstep groups examples by prefix length: 52 early-step examples use prefixes covering steps 1–2, 34 mid-step examples use prefixes covering steps 1–4, and 33 late-step examples use prefixes covering steps 1–6.
+Each example requires translating the next natural-language instruction given the corresponding NL–DSL history.
+Project-Level Translation
+In the project-level setting, the model is provided with the complete crochet instruction in natural language together with the corresponding product image, and must generate an entire CrochetPARADE program.
+This variant is globally self-contained but considerably more challenging than the step-level task: models must track stitch states over long horizons, resolve ambiguities in natural language, and produce code that is both syntactically valid and semantically aligned with the final design.
+This setting reflects how crochet instructions are used in practice, where each step depends on the correctness of all preceding steps.
+Image grounding is especially helpful for interpreting repeated motifs, symmetry, shaping, and termination conditions that may be under-specified in text alone.
+Evaluation.
+Because crochet patterns are inherently free-form (where multiple distinct programs can yield the same final product and a single natural-language instruction may admit several semantically equivalent DSL realizations), there is no canonical gold program for Task D.
+Exact string matching would therefore misjudge many correct solutions.
+Instead, CrochetBench evaluates correctness through functional executability using the CrochetPARADE validator, which checks whether a predicted DSL program is syntactically valid, structurally consistent, and fully executable.
+We use two complementary evaluation settings.
+For step-level translation, we report the Valid Pattern Rate, defined as the proportion of generated DSL steps that successfully compile.
+For project-level translation, we compute the Valid Pattern Rate for full programs and, for those that compile, render the executable portion into a crochet-like image and compute its DINO Similarity (Oquab et al., 2023) to the ground-truth product image, providing a coarse measure of semantic fidelity beyond syntax.
+To diagnose failure modes, we further identify the first point of failure for each invalid prediction and categorize it using our fine-grained error taxonomy (Appendix E), enabling us to distinguish local symbolic errors from broader state-tracking failures or misinterpretations of the natural-language instruction.
+Model Size Stitch Recognition (%) Instr. Sel. (%) Instr. Gen. (%)
+Prec Rec F1 Acc BLEU R-L ChrF
+Open Source BLIP-2 Flan-T5 XL 3B 29.53 23.03 22.50 25.62 0.22 9.30 9.45
+Google Gemma 3 4B 20.54 10.21 12.65 24.94 0.11 3.67 5.59
+Google Gemma 3 27B 17.19 18.14 16.05 24.94 0.40 5.17 6.55
+DeepSeek-VL 7B 54.47 74.76 60.60 28.92 1.30 19.39 18.14
+Qwen2-VL 7B 54.14 69.74 58.16 41.96 1.67 21.10 15.99
+Qwen2-VL 72B 71.86 42.68 50.19 68.85 2.25 21.43 19.82
+Closed Source GPT-4o 62.14 59.39 58.01 58.11 3.38 23.57 24.07
+Gemini 2.5 Flash-Lite 74.49 49.77 56.83 55.63 4.93 25.92 30.50
+Claude Sonnet 4 78.61 53.12 60.94 57.39 3.35 25.12 23.16
+Finetuned Qwen2-VL (finetuned) 7B - - - - 5.64 $\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}(\uparrow+3.97)$} 25.10 $\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}(\uparrow+4.00)$} 22.39 $\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}(\uparrow+6.40)$}
+Table 2: Combined evaluation results across all three CrochetBench tasks: Stitch Recognition, Instruction Selection, and Instruction Generation.
+Best values are bold; second-best are underlined.
+For the finetuned Qwen2-VL model, values in parentheses indicate absolute improvements over the pretrained Qwen2-VL 7B baseline.
+...A Midnight Blue ... ...B Copper Brown ... Ch4. Join with sl st to first ch to form ring ... 4th rnd: Sl st in next dc. Ch 2. 3 hdc in same dc as sl st. *Bobble in next st... Tassels (make 7)...
+(a) Ground Truth
+... With Color A (Blue), make a magic ring ... Join Color B (Brown) ... Rnd 4:Sl st into ch-2 sp. Ch 1, (Bobble ... Make 7 tassels with Color B and attach one to each ch-2 point of the star...
+(b) Gemini
+... Star-Shaped Crochet Blanket Pattern ... Yarn: Brown and Blue... With Brown yarn, make a magic ring. ... Tassels: Cut 4 lengths of Brown yarn ...
+(c) GPT-4o
+... 8-Point Star Afghan ... Worsted weight yarn in blue and brown ... With blue, ch 4, join with sl st to form ring. ... Tassels: Cut ...
+(d) Claude
+... Star Motif Crochet Pattern ... Worsted weight yarn in navy blue and brown ... Tassels... Round 1:Ch 10, join with sl st to form a ring ...
+(e) Qwen-VL-7B
+ch 100, sc in 2nd ch from hook, sc in each ch across, join with sl st in 1st sc, 100 loops made.
+(f) DeepSeek VL
+Figure 2: Case study for Task C: Instruction Generation.
+Each row shows the DSL-rendered output generated from the model’s natural-language instructions and the color-coded instruction extract below it.
+Matching colors denote semantically corresponding elements across the reference and model outputs, while red marks incorrect or hallucinated steps.
+The ground truth is a seven-point star with alternating blue and brown yarn and tassels attached at each point.
+Gemini and GPT-4o generate structured and mostly coherent instructions but misconstruct the global geometry, producing a circular motif rather than a star.
+Claude and Qwen2-VL-7B misinterpret the shape more severely, producing circular or distorted wireframe-like forms.
+DeepSeek-VL collapses entirely into a degenerate single-loop pattern.
+Gemini is the only model to explicitly recognize the motif as a seven-point star, but its instructions still fail to produce the correct star topology.
+To ensure transparency, we include the fully generated instructions for all models in Appendix F.
+4 Experiments
+We evaluate a representative set of widely used vision–language models spanning open and closed ecosystems.
+For open source models, we include BLIP-2 Flan-T5 XL (Li et al., 2023), Google Gemma 3 (4B and 27B) (Team et al., 2024), DeepSeek-VL 7B (Lu et al., 2024), and Qwen2-VL (7B and 72B) (Wang et al., 2024), covering a range of architectures and parameter scales.
+For closed source models, we evaluate GPT-4o (Hurst et al., 2024) , Gemini 2.5 Flash-Lite (Comanici et al., 2025), and Claude Sonnet 4 (Anthropic, 2025), which represent the strongest publicly accessible multimodal systems.
+These models span diverse architectures and parameter scales, providing a diverse and meaningful basis for assessing current multimodal capabilities on perception, retrieval, and procedural reasoning tasks.
+We evaluate all models on the four CrochetBench tasks described in Section 3, using the same metrics and evaluation protocols throughout.
+The prompts used for evaluation are provided in Appendix C.
+Results and analysis are listed below.
+Perception and grounding improve with scale, but procedural generation collapses.
+Table 2 summarizes results across Stitch Recognition (Task A), Instruction Selection (Task B), and Instruction Generation (Task C).
+Closed-source models achieve the strongest recognition performance, with Claude Sonnet 4 obtaining the highest F1 score (60.94%), and DeepSeek-VL 7B leading among open models (60.60%).
+A clear precision–recall trade-off emerges across model scales: smaller models such as DeepSeek-VL 7B and Qwen2-VL 7B achieve higher recall but lower precision, whereas larger models such as Qwen2-VL 72B and closed-source models favor precision over recall.
+This pattern reflects different grounding strategies: smaller models tend to broadly enumerate common stitch types to maximize coverage, while larger models adopt a more selective, conservative approach, yielding higher precision but incomplete visual grounding.
+Although larger models capture more fine-grained visual cues, accuracy remains far from saturated, and Instruction Selection shows similarly limited progress, with most large and closed-source models clustering around 55–60% accuracy, indicating that visual–textual alignment still depends on shallow correlations rather than robust grounding.
+These limitations become more pronounced in Task C.
+Natural-language instruction generation remains extremely challenging for all model, with BLEU, ROUGE-L, and ChrF scores uniformly low; even the strongest system, Gemini 2.5 Flash-Lite, achieves only 4.93% BLEU and 30.50% ChrF.
+The sharp drop from Tasks A and B to C shows that models capable of recognizing stitches or retrieving plausible text still fail to synthesize coherent multi-step procedures, reflecting fundamental gaps in procedural reasoning, symbolic consistency, and pattern-structure understanding.
+Surface-level fluency does not imply procedural correctness.
+Figure 2 presents a case study comparing model-generated natural-language instructions with their corresponding DSL renderings.
+Qwen2-VL-7B and DeepSeek-VL collapse into non-star geometries, revealing unstable procedural logic.
+GPT-4o and Claude produce structurally well-formed, human-like crochet instructions, including explicit materials sections and round-level organization, and correctly capture local yarn colors.
+However, they fundamentally misinterpret the global motif: GPT-4o reconstructs a four-point star and begins with the brown yarn instead of blue, while Claude generates an eight-point motif rather than the intended seven.
+Gemini most accurately identifies the seven-point structure and selects plausible constructs such as bobbles for the star tips, but structural inconsistencies remain and yield visibly distorted shapes.
+These examples demonstrate that models can generate fluent, crochet-like descriptions while failing to preserve the algorithmic structure required for faithful pattern synthesis.
+Supervised finetuning improves fluency but not procedural correctness.
+We evaluate supervised finetuning on CrochetBench-C using an 80/5/15 stratified train/validation/test split with full-parameter updates.
+As shown in Table 3 (bottom line), finetuning substantially improves all generation metrics: BLEU increases from 1.67% to 5.64% (+238%), ROUGE-L from 21.10% to 25.10% (+19%), and ChrF from 15.99% to 22.39% (+40%).
+The finetuned 7B model outperforms all evaluated closed-source systems, including GPT-4o, Claude Sonnet 4, and Gemini 2.5 Flash-Lite.
+Despite these gains, absolute performance remains low (5.64% BLEU), indicating that finetuning does not resolve core procedural failures.
+Qualitative analysis shows that improvements are largely superficial: outputs exhibit more consistent formatting, improved stitch vocabulary, and better adherence to crochet conventions, but still contain frequent stitch-count errors, malformed repeats, and incorrect global structure (e.g., rounds vs.
+rows, increases/decreases).
+These findings suggest that finetuning primarily enhances surface fluency rather than inducing robust algorithmic or structural reasoning.
+Figure 3: Task D step-level translation results showing the proportion of generated DSL lines that successfully compile for early (Steps 1–2), middle (Steps 1–4), and late (Steps 1–6) stages of crochet patterns.
+Across all models, valid pattern rates increase as more context is provided, but overall accuracy remains low.
+Even the strongest models struggle in early steps, indicating difficulty establishing correct stitch state and structural dependencies.
+Larger models (e.g., Qwen2-VL-72B and Gemma-27B) do not consistently outperform their smaller counterparts, highlighting scale alone does not improve program-level structural reasoning.
+Early-step instability reveals limits of procedural reasoning.
+Figure 3 shows step-level results on Task D.
+Valid Pattern Rate increases with pattern depth but remains low overall: most models achieve under 15% validity in the first two steps, improve modestly in steps 3–4, and reach only 55–65% in later steps.
+This pattern reflects the difficulty of the initial steps, which must correctly initialize the program state such as defining stitch variables and maintaining balanced grouping.
+Errors made early propagate irreversibly, and later correctness often occurs only when the initial state is accidentally valid, indicating reliance on continuation heuristics rather than genuine procedural understanding.
+Larger models do not consistently perform better: Qwen2-VL-72B underperforms Qwen2-VL-7B, and Gemma-27B underperforms Gemma-4B, suggesting that increased capacity improves descriptive fluency more readily than symbolic stability, and that scaling alone is insufficient for grammar-sensitive procedural tasks.
+Figure 4: Distribution of project-level DSL translation outcomes for each model, broken down into valid outputs and error categories.
+Across all models, invalid programs dominate, with most failures arising from undefined stitches, unbalanced brackets, and multiple-reference errors.
+The spread of error types further illustrates the difficulty of maintaining global consistency and symbolic correctness when generating full crochet programs.
+Project-level DSL synthesis reveals severe long-range structural failures, and does not improve with model scale.
+Figure 4 highlights the fragility of full-program CrochetPARADE synthesis.
+Valid outputs are rare: even the strongest models (Claude, Gemini, and Qwen2-VL-7B) produce only 5–8% executable programs, while most others fall below 3%.
+The most frequent failure modes—undefined stitches and unbalanced brackets—indicate poor control over the DSL vocabulary and grouping structure.
+Additional errors, including multiple references, non-adjacent labels, and runtime failures, further reveal difficulties in maintaining consistent state and long-range structural dependencies across an entire pattern.
+Interestingly, scaling does not alleviate these issues and can exacerbate them.
+Larger models exhibit substantially higher rates of undefined-stitch errors, suggesting a tendency toward uncontrolled symbolic invention (e.g., Qwen2-VL-72B: 72.0% vs.
+Qwen2-VL-7B: 13.6%; Gemma-3-27B: 42.6% vs.
+Gemma-3-4B: 27.4%).
+Image-based similarity confirms lack of global structural fidelity.
+Compilation verifies syntactic and structural correctness but cannot determine whether two DSL programs are semantically equivalent.
+To address this gap, we compute DINO similarity between the target crochet product image and the rendering produced from each model’s executable program (valid outputs only).
+Figure 5 shows that similarity scores remain uniformly low across all models (0.10–0.17), far below the typical threshold for visually matched crochet images.
+Even when a model produces a compilable DSL program, the resulting rendering generally bears little resemblance to the intended pattern, indicating that syntactic validity does not imply correct procedural structure.
+The consistently low similarities reinforce that current multimodal LLMs fail to capture the global geometry and layout required for visually faithful crochet synthesis.
+Figure 5: Task D Project-level translation results evaluating with DINO similarity between ground-truth images and DSL-rendered outputs generated from each model’s DSL program (valid executable portion only).
+The red line marks an approximate “good” similarity threshold.
+All models fall well below this level, indicating that even executable DSL programs rarely reproduce the correct visual structure of the intended crochet design.
+5 Related Work
+Multimodal benchmarks have traditionally emphasized descriptive image–text alignment, as in COCO (Lin et al., 2014) and Flickr30k (Plummer et al., 2015).
+More recent datasets extend to instructional or procedural understanding, including Recipe1M+ (Marin et al., 2018) and large-scale instructional video corpora such as YouCook2 and HowTo100M (Zhou et al., 2018; Miech et al., 2019).
+However, these benchmarks primarily evaluate semantic alignment, retrieval, or temporal recognition, and do not assess whether a model can generate or execute a correct procedure.
+Texile crafts provide precisely such a domain.
+Crochet patterns specify symbolic, stepwise procedures that determine the topology and geometry of a final physical artifact.
+Prior work in this area including Digital Crochet (Seitz et al., 2022) and Neural Inverse Knitting (Kaspar et al., 2019) demonstrates the feasibility of representing textile structures in machine-readable form but remains limited in scale and modality.
+By moving from general instructional data to a structured craft domain, we enable multimodal models to be evaluated on artifact-centric procedural reasoning rather than temporal action recognition.
+CrochetBench builds on this emerging direction by providing thousands of real crochet patterns with paired images and natural-language instructions.
+To evaluate procedural correctness, CrochetBench adopts an executable domain-specific language (CrochetPARADE), linking our tasks to program synthesis benchmarks such as HumanEval (Chen, 2021b), MBPP (Austin et al., 2021), and Spider (Yu et al., 2018).
+In multimodal settings, image-to-program benchmarks such as Im2LaTeX-100K (Deng et al., 2017) and pix2code (Beltramelli, 2018) similarly leverage executable formalisms for rendering-based evaluation.
+CrochetBench extends this executable perspective to textile crafts: patterns compile to structured instructions that can be rendered and validated, providing functional evaluation that tests whether a model’s output actually works.
+This offers a lightweight alternative to domains such as chemistry or cooking, where validating a procedure requires physical or chemical experiments that are slow, costly, or impractical to scale.
+6 Conclusion
+CrochetBench provides a structured benchmark for assessing whether multimodal LLMs can move from recognizing visual content to executing the step-by-step procedures required to produce a crochet pattern.
+Across all four tasks, models demonstrate a consistent gap: they can identify stitches and retrieve plausible instructions, but they fail to generate structurally valid procedures or produce executable programs.
+Qualitative analysis reveals that fluent outputs often misrepresent global structure, and scaling further exacerbates execution failures due to uncontrolled invention of stitch types.
+These results highlight the need for methods that target structured, long-horizon reasoning.
+Limitations
+While CrochetBench enables executable evaluation of visual procedural reasoning, it has several limitations.
+First, the benchmark focuses on a single creative domain, crochet, which although representative of structured symbolic procedures, may not capture all forms of procedural reasoning found in domains such as robotics, mechanical assembly, or scientific experimentation.
+As a result, generalization to other domains should be interpreted with caution.
+Second, executable correctness is assessed through the CrochetPARADE domain specific language, which necessarily abstracts certain aspects of real world crochet practice.
+Multiple distinct procedures can produce visually similar artifacts, and some valid human written patterns may be penalized due to constraints of the language or limitations of the rendering process.
+Although we mitigate this issue using functional validation and image based similarity metrics, these signals remain imperfect proxies for true semantic equivalence.
+Finally, while we evaluate a diverse set of state of the art vision language models, our finetuning experiments are limited to supervised learning on a single architecture.
+Exploring alternative training paradigms, such as program guided learning or learning with execution feedback, remains an important direction for future work.
+Ethical Considerations
+We acknowledge that the original crochet pattern PDFs are protected under copyright and therefore do not distribute raw files or full texts.
+Instead, we release only structured JSON annotations generated with GPT, reference URLs to the original sources, and our parsing and annotation scripts.
+The benchmark is provided strictly for non-commercial academic use.
+This approach enables reproducible research while respecting intellectual property and ensuring that our dataset serves as a tool for studying structured generation rather than redistributing creative works.
+Appendix A Crochet Primer
+Crochet patterns describe how to construct a textile artifact through a sequence of symbolic stitch instructions.
+Each instruction specifies an operation performed with a hook and yarn, and the resulting pattern is defined by the order, repetition, and spatial arrangement of these stitches.
+This appendix summarizes only the conventions needed to interpret the examples in our benchmark.
+Basic stitch types.
+Crochet relies on a small vocabulary of atomic stitches, each producing a loop with a characteristic height and structure. The most common stitches in U.S. notation are:
+  • ch (chain): foundational stitch used to begin rows or rounds.
+  • sc (single crochet): a short, dense stitch.
+  • hdc (half double crochet) and dc (double crochet): taller stitches that build height more quickly.
+  • sl st (slip stitch): a joining stitch used for connecting motifs or closing rounds.
+  • bobble (bobble stitch): a cluster of 5 partially completed double crochet stitches closed together into a single stitch.
+These stitches can be combined in rows (worked back and forth) or rounds (worked in a circle).
+Pattern syntax and structure.
+Crochet instructions follow a compact symbolic notation. A pattern is organized into rows or rounds, each specifying a sequence of stitches. For example:
+Row 3: Ch 1, sc in each st across, turn.
+Instructions may include:
+  • Repetition: indicated by parentheses and a multiplier, e.g., (sc, ch 1) 3 times.
+  • Increases/decreases: e.g., 2 sc in next st (increase) or sc2tog (single-crochet two stitches together; decrease).
+  • Stitch counts: patterns often end rows or rounds with “—N sc,” indicating the number of stitches that should remain.
+Relationship to symbolic representations.
+Each crochet instruction corresponds to a local modification of the fabric’s topology.
+This makes crochet patterns naturally suited to symbolic or program-like representations such as CrochetPARADE, which encode stitches as structured primitives with explicit control flow (loops, groups, labels).
+Because stitch sequences fully determine the geometry of the final artifact, correctness can be assessed by verifying the structure of the generated program or by rendering the corresponding stitch graph.
+This primer covers the minimal terminology required to interpret our dataset and evaluation tasks.
+For readers interested in additional background, standard crochet references provide extended stitch catalogs and diagram conventions.
+Appendix B Additional Dataset Statistics
+CrochetBench supports diverse real-world crochet practices, with project types ranging from simple accessories to complex garments.
 Figure 6 lists the ten most common categories by frequency.
 The majority of patterns belong to a small number of dominant types—Afghans and Blankets alone account for over one-quarter of the dataset.
-More details can be found at Appendix B.
-Each pattern is labeled with one of four primary skill levels—beginner, easy, intermediate, or experienced.
+Each pattern is labeled with one of four primary skill levels, including beginner, easy, intermediate, or experienced.
 This allows for stratified evaluation across complexity tiers.
 Figure 7 shows the skill level distribution, which is strongly skewed toward beginner-friendly content.
-Only one pattern (0.02%) is missing a skill level label.
 More details can be found at Appendix B.
@@
 This lexical mapping enables tasks such as vocabulary translation, sequence generation, and instruction validation.
-Beyond raw instructions, the structured schema also records rich metadata, including gauge, hook size, yarn weight, and measurements.
-A representative dataset entry is provided in Appendix 10.
-Overall, CrochetBench provides a rich resource for multimodal modeling, symbolic reasoning, and structure-aware generation.
-Its coverage across diverse categories and complexity levels enables broad benchmarking of both open-ended generation and instruction fidelity tasks.
-Table 1: Overall statistics of the CrochetBench dataset.
-Total Patterns Image Coverage Avg. Instr. Length #Project Types
-CrochetBench 6,085 98.77% 3,216 characters 55
-3 Experiment
-We empirically evaluate CrochetBench by defining a set of structured tasks and benchmarking a diverse pool of multimodal large language models.
-Our experiments are designed to test models across progressively challenging stages, from low-level recognition to high-level executable synthesis.
-Benchmark Tasks.
-Table 2 summarizes the four evaluation tasks in CrochetBench, which progress systematically from recognition to comprehension, generation, and ultimately executable synthesis.
-Task A (Stitch Recognition) evaluates a model’s ability to detect symbolic primitives in crochet images, establishing the foundation for multimodal perception.
-Task B (Instruction Selection) requires models to align visual evidence with candidate textual instructions, thereby testing multimodal grounding and fine-grained comprehension.
-Unlike conventional description tasks, the candidates are procedural steps rather than captions, and correct selection often requires reasoning about how local steps contribute to the final product.
-Task C (Instruction Generation) advances from comprehension to open-ended production, challenging models to generate natural-language procedural instructions that are both perceptually grounded and linguistically faithful to domain conventions.
-Evaluation here emphasizes lexical and symbolic fidelity, but does not directly capture structural validity.
-Task D (Instruction-to-DSL Translation) addresses this gap by requiring models to output a compilable program in a domain-specific language (DSL).
-The step-level variant tests local semantic grounding, while the project-level variant demands global structural consistency across the entire pattern.
-Compilation-based evaluation directly measures executable faithfulness, ensuring that generated instructions are not only linguistically plausible but also structurally sound.
-Together, these tasks form a structured evaluation ladder that moves from perceptual recognition to programmatic execution, probing both low-level perception and high-level symbolic reasoning.
-Table 2: Overview of benchmark tasks in CrochetBench. Tasks progress from recognition to comprehension, generation, and executable synthesis.
-ID Ability Tested Task Evaluation Metrics Test Size
-A Recognition Stitch Recognition F1, Precision, Recall
-6,009
-(CrochetBench-A)
-B Comprehension Instruction Selection Accuracy
-6,003
-(CrochetBench-B)
-C Generation Instruction Generation BLEU, ROUGE, ChrF
-6,009
-(CrochetBench-C)
-D Formalization Instr.-to-DSL (Step) Valid Pattern Rate
-119
-(CrochetBench-Dstep)
-Instr.-to-DSL (Project) Valid Pattern Rate
-100
-(CrochetBench-Dproj)
-Model Selection.
-We evaluate a diverse set of vision–language models (VLMs) spanning both open-source and closed-source families.
-The open-source group includes Salesforce BLIP-2 Flan-T5 XL (3B), a perception-focused baseline widely used in image–text tasks; Google Gemma 3 (4B) and Qwen2-VL (7B), two recent models trained with large-scale multimodal alignment; and DeepSeek-VL (7B), a larger open-source model designed for stronger vision–language reasoning.
-On the closed-source side, we assess GPT-4o, Gemini 2.5 Flash-Lite, and Claude Sonnet 4, which represent state-of-the-art commercial VLMs that push the limits of multimodal reasoning.
-This selection spans perception-heavy baselines to general-purpose multimodal reasoning systems, enabling us to evaluate capabilities across both research-grade and production-grade settings.
-[^2] We use the term “VLM” broadly to include both traditional vision–language models (e.g., BLIP-2) and modern multimodal large language models (e.g., GPT-4o, Gemini, Claude).
-3.1 Task A: Stitch Recognition
-The first task evaluates a model’s ability to identify crochet stitch types from an image of a finished product.
-We construct CrochetBench-A, a subset of 6,009 examples from the full benchmark, where each product image is paired with ground-truth stitch annotations.
-These labels are derived from the official pattern instructions and normalized into a standardized set of stitch abbreviations (e.g., sc, hdc, dc) to ensure consistency across patterns.
-Unlike standard image classification, this is a multi-label prediction problem: multiple stitches may co-occur within the same image, often with subtle visual differences in texture and geometry.
-This task therefore probes fine-grained visual grounding of structured crochet semantics.
-Evaluation.
-For each example, we compute overlap between the predicted and reference stitch sets.
-True Positives (TP) are stitches correctly predicted; False Positives (FP) are stitches predicted but not in the reference; and False Negatives (FN) are stitches in the reference but missed by the model.
-From these counts, we compute precision (fraction of correct predictions among all predictions), recall (fraction of ground-truth stitches recovered), and F1 score (harmonic mean).
-Metrics are averaged across examples to provide overall performance.
-This formulation rewards models that recover all present stitches while avoiding spurious predictions.
-Accurate stitch recognition is foundational for the benchmark, as later tasks (e.g., instruction selection and instruction generation) depend on robust detection of stitch primitives.
-Table 3: Evaluation results on the Stitch Recognition task. We report Precision, Recall, and F1. Best results are bold; second-best are underlined.
-Model Size Precision Recall F1
-Open Source Salesforce BLIP-2 Flan-T5 XL 3B 0.2953 0.2303 0.2250
-Google Gemma 3 4B 0.2054 0.1021 0.1265
-DeepSeek-VL 7B 0.5447 0.7476 0.6060
-Qwen2-VL 7B 0.5414 0.6974 0.5816
-Closed Source GPT-4o 0.6214 0.5939 0.5801
-Gemini 2.5 Flash-Lite 0.7449 0.4977 0.5683
-Claude Sonnet 4 0.7861 0.5312 0.6094
-Results.
-Claude Sonnet 4 achieves the best overall F1 score, demonstrating strong precision in stitch recognition.
-Among open-source models, DeepSeek-VL performs best, with notably high recall, while Qwen2-VL offers a competitive balance.
-These results highlight the gap between commercial VLMs and open-source alternatives, but also suggest that large open-source models are beginning to approach closed-source performance in fine-grained recognition.
-3.2 Task B: Instruction Selection
-The second task evaluates whether a model can correctly align a finished crochet product image with its corresponding natural-language instruction.
-We construct CrochetBench-B, a subset of 6,003 examples, where each item consists of one correct instruction and three distractor instructions sampled from the same project category (e.g., hats, rugs).
-Distractors are carefully chosen to share structural motifs and vocabulary, thereby increasing difficulty and requiring fine-grained visual–textual alignment rather than reliance on superficial cues.
-The answer distribution across options is approximately uniform (A: 24.9%, B: 25.7%, C: 23.7%, D: 25.7%).
-Evaluation.
-To enable scalable assessment, we formulate the task as a 4-way multiple-choice question (MCQ).
-The model must select one option (A–D), with exactly one correct answer.
-Predictions are extracted using a lightweight regex-based method that identifies explicit letter outputs (e.g., “A”, “The answer is B”).
-If no parsable choice is found, the response is marked as unanswered.
-Accuracy is reported as the evaluation metric.
-This task provides a controlled measure of visual grounding and semantic alignment between image content and procedural text, without requiring generative modeling.
-It highlights the challenge of distinguishing between subtle visual cues and domain-specific terminology, which is essential for bridging perception and structured instruction understanding.
-Table 4: Evaluation results on the Instruction Selection task (4-way multiple choice). We report Accuracy. Best results are bold; second-best are underlined.
-Model Size Accuracy
-Open Source Salesforce BLIP-2 Flan-T5 XL 3B 0.2562
-Google Gemma 3 4B 0.2494
-DeepSeek-VL 7B 0.2892
-Qwen2-VL 7B 0.4196
-Closed Source GPT-4o 0.5811
-Gemini 2.5 Flash-Lite 0.5563
-Claude Sonnet 4 0.5739
-Results.
-GPT-4o achieves the highest accuracy, closely followed by Claude and Gemini, all of which substantially outperform open-source VLMs.
-Among open-source models, Qwen2-VL is the strongest, while BLIP-2 and Gemma remain near chance level, underscoring the difficulty of visually grounded instruction matching in this domain.
-3.3 Task C: Instruction Generation
-The third task evaluates a model’s ability to generate natural-language crochet instructions directly from an image of a finished product.
-We construct CrochetBench-C, a subset of 6,009 examples, where each product image is paired with its full pattern.
-Unlike classification or selection, this task requires producing a multi-step sequence that follows domain-specific syntax, ordering, and stitch logic.
-Outputs are expected to resemble real-world crochet patterns, written line by line (e.g., “Rnd 1: ch 4, 6 sc in ring”).
-This setting challenges models to translate visual evidence into coherent procedural text that preserves both symbolic accuracy and structural consistency.
-Evaluation.
-We evaluate generation quality using standard text-generation metrics and a domain-specific structural measure.
-BLEU and ROUGE-L capture lexical and n-gram overlap with the reference instructions.
-ChrF, computed over character n-grams, provides a finer-grained signal of similarity and is particularly suited to stitch abbreviations.
-While these metrics measure surface-level overlap, they do not directly test whether the generated instructions are structurally consistent (e.g., balanced stitch counts, valid round progression).
-To address this limitation, we complement Task C with Task D, where compilation-based evaluation of DSL translations directly assesses executable consistency.
-Together, these tasks allow us to probe both linguistic fidelity and procedural correctness.
-Table 5: Evaluation results on the Instruction Generation task. We report BLEU, ROUGE-L, and ChrF. Higher is better. Best results are bold; second-best are underlined.
-Model Size BLEU ROUGE-L ChrF
-Open Source Salesforce BLIP-2 Flan-T5 XL 3B 0.0021 0.0926 9.32
-Google Gemma 3 4B 0.0010 0.0329 5.17
-Qwen2-VL 7B 0.0160 0.2084 15.76
-DeepSeek-VL 7B 0.0133 0.1968 18.12
-Closed Source GPT-4o 0.0333 0.2353 23.80
-Gemini 2.5 Flash-Lite 0.0482 0.2583 30.20
-Claude Sonnet 4 0.0331 0.2516 22.95
-Results.
-Gemini achieves the strongest overall performance across all metrics, substantially outperforming other closed-source models.
-GPT-4o and Claude follow closely, while open-source models lag behind with significantly lower BLEU and ChrF scores.
-This gap highlights the difficulty of generating structurally faithful crochet instructions, which requires models to capture both visual details and domain-specific procedural logic.
-3.4 Task D: Instruction-to-DSL Translation
-We construct two subsets for Task D: CrochetBench-Dstep (119 items) and CrochetBench-Dproj (100 items).
-For CrochetBench-Dstep, we manually annotate the correct CrochetPARADE DSL for the previous context of each pattern (a prefix of NL (natural language) –DSL pairs).
-The model is then given this prefix along with the next natural-language instruction, and must predict the corresponding DSL line.
-Predictions are evaluated by checking compilation validity with the CrochetPARADE validator.
-For CrochetBench-Dproj, we adopt a few-shot prompting setup rather than full manual annotation.
-Annotators provide a single reference program in CrochetPARADE as context, which is paired with a new natural-language instruction and its product image.
-The model must then generate a complete DSL program, and outputs are assessed by whether they compile fully or partially using the validator.
-Design Rationale.
-Crochet patterns admit many valid DSL realizations (e.g., alternative groupings or equivalent constructs), meaning that no single gold reference is canonical.
-A reference-based metric could unfairly penalize models that produce semantically correct but structurally different programs.
-By instead relying on the validator, CrochetBench shifts evaluation away from surface-level string matching toward functional executability—the central criterion for crochet synthesis.
-This design is consistent with other program synthesis benchmarks, such as semantic parsing and SQL generation, where execution accuracy is often preferred over exact string match.
-3.4.1 Step-Level Translation
-In the step-level setting, the model receives a prefix of NL–DSL pairs and must generate the DSL line corresponding to the next natural-language instruction.
-This setup reflects an incremental synthesis process in which correctness depends on maintaining stitch-level consistency across steps.
-Since crochet patterns are inherently stateful, earlier context is critical for resolving constructs such as increases, repeats, and turning chains.
-To capture progression through a pattern, we sample 52 early examples (steps 1–2), 34 mid examples (steps 3–4), and 33 late examples (steps 5–6).
-Step-level inputs are formatted as:
-$\text{Prefix (NL--DSL pairs)}+\text{Next NL instruction}\;\;\rightarrow\;\;\text{Next DSL line}.$
-This formulation allows models to generate locally plausible DSL steps, but global correctness ultimately requires stronger contextual reasoning.
-Evaluation.
-We evaluate models using Compilation Success Rate (CSR), defined as the proportion of generated DSL outputs that compile successfully with the CrochetPARADE validator.
-This validator-based metric is a key strength of CrochetBench: unlike reference-based string matching, CSR directly measures whether generated programs are executable, rewarding functional correctness rather than surface similarity.
-To better understand systematic errors, we also conduct fine-grained error analysis across four categories: (1) syntax structure errors, (2) stitch definition errors, (3) labeling and reference errors, and (4) structural or formatting issues.
-A detailed taxonomy with illustrative examples is provided in the Appendix.
-CrochetBench further supports multimodal verification, as CrochetPARADE programs can be rendered into simulated product images.
-These renderings can be compared with ground-truth product images using pretrained vision–language models (e.g., CLIP), enabling evaluation of both structural validity and visual faithfulness.
-While current model performance leaves ample room for progress, the combination of validator-based execution checks and prospective image-grounded verification establishes CrochetBench as a uniquely rigorous testbed for structured multimodal reasoning.
-Table 6: Verification results across models, grouped by category.
-Model CSR (%) Undef. Br. MRef Other Tot. Err.
-Open Source
-Salesforce BLIP-2 Flan-T5 XL 4.2 29.8 48.2 20.2 1.8 114
-Google Gemma 3 3.4 26.1 10.4 63.5 115
-DeepSeek-VL 32.8 36.2 38.8 11.2 13.8 80
-Qwen2-VL 35.3 42.9 39.0 1.3 16.8 77
-Closed Source
-GPT-4o 36.1 39.5 43.4 17.1 76
-Gemini 2.5 Flash-Lite 41.2 51.4 21.4 5.7 21.5 70
-Claude Sonnet 4 52.1 45.6 28.1 26.3 57
-Results.
-Step-level translation is locally tractable: models often produce syntactically valid lines, but compilation failures reveal semantic underspecification.
-Claude achieves the highest valid rate (52.1%), while DeepSeek-VL and Qwen2-VL are the strongest among open-source models.
-Errors are dominated by syntax (brackets, references) and undefined stitches, underscoring the difficulty of maintaining state consistency even when local syntax is correct.
-3.4.2 Project-Level Translation
-In the project-level setting, the model is provided with the complete crochet instruction in natural language together with the corresponding product image, and must generate an entire CrochetPARADE program.
-This variant is globally self-contained but considerably more challenging than the step-level task: models must track stitch states over long horizons, resolve ambiguities in natural language, and produce code that is both syntactically valid and semantically aligned with the final design.
-Image grounding plays a crucial role in disambiguating constructs such as repeated motifs, symmetry, and termination conditions.
-Evaluation.
-We assess model outputs using two complementary metrics.
-The first is Compilation Success Rate (CSR), and the second is Partial Executable Rate (PER), which measures the average fraction of a program that compiles successfully before failure.
-While CSR captures all-or-nothing executability, PER provides a finer-grained view of structural alignment, offering credit to models that generate correct prefixes even if the full program does not compile.
-Error types are categorized using the same taxonomy as in the step-level evaluation.
-Table 7: Project-level verification results. We report Compilation Success Rate (CSR) and Partial Executable Rate (PER). Best scores are bold, second-best are underlined.
-Model Size CSR (%) PER (%)
-Open Source Salesforce BLIP-2 Flan-T5 XL 3B 1.0 0.00
-Google Gemma 3 4B 1.6 5.29
-DeepSeek-VL 7B 8.1 37.49
-Qwen2-VL 7B 21.0 30.28
-Closed Source GPT-4o 4.0 2.76
-Gemini 2.5 Flash-Lite 4.0 5.67
-Claude Sonnet 4 5.0 8.16
-Table 8: Project-level error type distribution across models (percentage).
-Abbreviations: Undef.
-(Undefined), Br.
-(Brackets), Lbl.
-(Labels Missing), Non-adj.
-(Non-adjacent), MRef (Multi-Reference), Other (Other/Runtime).
-Model Undef. Br. Lbl. Non-adj. MRef Other
-Open Source
-Salesforce BLIP-2 Flan-T5 XL 2.0 37.4 58.6 2.0
-Google Gemma 3 39.3 14.8 11.5 8.2 14.8 11.5
-DeepSeek-VL 36.3 45.1 4.4 11.0 3.3
-Qwen2-VL 25.0 12.5 7.8 51.6 3.1
-Closed Source
-GPT-4o 61.5 7.3 16.7 4.2 4.2 6.2
-Gemini 2.5 Flash-Lite 51.1 18.1 9.6 8.5 4.3 8.6
-Claude Sonnet 4 46.3 14.7 21.1 7.4 5.3 5.3
-Results.
-Project-level translation remains highly challenging.
-Qwen2-VL achieves the best valid rate (21.0%) and strong PER, surpassing all closed-source systems.
-DeepSeek-VL demonstrates robust partial executability despite a lower valid rate.
-By contrast, GPT-4o, Gemini, and Claude achieve lower scores, highlighting that even state-of-the-art VLMs struggle with long-range structural consistency in executable synthesis.
-Error analysis (Table 8) reveals that closed-source models often fail due to label and reference inconsistencies, while open-source models more frequently exhibit undefined stitches and syntax errors.
-4 Discussion
-The results across CrochetBench highlight both the promise and current limitations of Visual language models in bridging perception, procedural reasoning, and executable synthesis.
-By structuring tasks in a progressive ladder, we expose clear gradients of difficulty: while contemporary models demonstrate competence in low-level recognition and mid-level comprehension, their performance declines substantially when asked to generate or formalize instructions into compilable domain-specific programs.
-This pattern underscores both methodological bottlenecks in multimodal grounding and fundamental challenges in symbolic reasoning over long-horizon structures.
-In Task A, stitch recognition is feasible for both open- and closed-source models, though precision–recall tradeoffs differ.
-Claude Sonnet 4 prioritizes precision for a higher F1, while DeepSeek-VL favors recall via richer visual encoders.
-This reflects divergent inductive biases: commercial models better regularize spurious predictions, while open-source models overpredict to capture subtle textures.
-Crucially, this is a multi-label, fine-grained texture classification task with symbolic implications, distinguishing it from generic object recognition.
-Baselines like BLIP-2, designed for captioning, underperform due to insufficient symbolic grounding (blip; clip).
-Task B amplifies these challenges.
-Closed-source models, especially GPT-4o, outperform open models by a wide margin, underscoring the difficulty of aligning visual cues with procedural semantics.
-Distractor instructions are intentionally plausible, demanding reasoning over local–global coherence, not just token overlap (GVE; VQA; miech2019howto100mlearningtextvideoembedding).
-Qwen2-VL’s competitive performance suggests that scale and pretraining diversity help, but commercial systems benefit from stronger instruction tuning and better vision–language alignment.
-Task C exposes the steep drop in open-ended procedural generation.
-Gemini 2.5 leads across BLEU, ROUGE, and ChrF (papineni2002bleu; lin2004rouge; popovic2015chrf), reflecting fluency in structured text.
-Yet absolute scores are low, and outputs often fail to conform to crochet-specific syntax.
-Open models hallucinate frequently, lacking exposure to domain-aligned distributions.
-This highlights a broader issue: fluency in general language modeling does not imply competence in structured, domain-specific generation (miech2019howto100mlearningtextvideoembedding; alayrac2022flamingovisuallanguagemodel).
-Task D proves most demanding.
-Even top closed-source models rarely exceed 6% project-level compilation success, while Qwen2-VL unexpectedly achieves 21%, suggesting better symbolic generalization under execution constraints.
-Syntax errors (e.g., undefined stitches, malformed brackets) dominate in open models, while closed models produce syntactically valid but semantically inconsistent programs.
-These complementary failure modes hint at differing generalization paths: symbolic robustness versus lexical fluency.
-Notably, Task D also tests 3D-aware procedural fidelity—models must translate visual or textual cues into symbolic programs that unfold into spatially coherent structures, not just grammatically valid sequences.
-Taken together, these results demonstrate that success in linguistic generation does not translate directly to executable synthesis.
-Performance decays sharply when evaluation moves from surface-level fidelity (BLEU, ROUGE) to structural validity (compilation), reinforcing the importance of execution-grounded metrics for procedural tasks.
-The relative strength of Qwen2-VL at the project level further suggests that progress may come not from scaling alone, but from architectural or training adjustments that better capture long-range dependencies and stateful operations.
-More broadly, CrochetBench exposes a critical frontier for multimodal reasoning: models must not only ground text in vision but also internalize procedural invariants that guarantee functional correctness.
-However, our analysis must be contextualized within certain constraints.
-The benchmarks reflect standardized stitch sets and normalized instructions, which, while ensuring comparability, simplify the variability encountered in real-world crochet practice (e.g., designer-specific shorthand, unconventional repeats).
-Moreover, evaluation metrics in Task C rely on string overlap, which may underestimate semantically correct but lexically divergent outputs.
-Even the compilation-based metric in Task D, though stronger, cannot assess visual fidelity unless paired with image-render verification.
-Finally, the sample size for project-level DSL translation remains limited, and absolute validity rates are low, constraining fine-grained statistical comparisons.
-5 Future Work
-A central avenue for future research is advancing models that translate free-form natural language crochet instructions into the formal CrochetPARADE DSL.
-As a domain-specific programming language, CrochetPARADE not only enables executable verification of multimodal outputs but also positions crochet as a program synthesis problem, where compilers map symbolic grammars into machine-executable instructions, extending ideas from prior DSL work in domains such as knitting (knitscript), graphics (ellis2018learninginfergraphicsprograms), and robotics (mu2024robocodexmultimodalcodegeneration).
-This perspective connects naturally to CAD/CAM integration: industrial crochet and warp-knitting machines (e.g., COMEZ, Jakob Müller) already rely on pipelines from graphical design interfaces, to intermediate graph-based representations, to low-level machine code.
-CrochetPARADE could serve as a standardized intermediate layer in this workflow, bridging human-facing authoring tools with machine-facing execution systems (text2cad).
-Several technical directions arise from this framing.
-First, incorporating explicit state-tracking mechanisms—whether through memory-augmented architectures (graves2016hybrid) or symbolic scaffolds (nye2021workscratchpadsintermediatecomputation)—could mitigate long-range inconsistencies in DSL translation.
-Second, multimodal pretraining enriched with procedural and topological domains (e.g., assembly instructions (cao2023instruction), instructional videos (miech2019howto100mlearningtextvideoembedding)) may narrow the gap between natural-language fluency and executable synthesis.
-Third, hybrid evaluation pipelines that combine compilation checks with visual render comparisons could more holistically assess structural and perceptual fidelity, building on metrics used in programmatic 3D generation (jain2022zero; paschalidou2021neuralpartslearningexpressive).
-Finally, CrochetBench offers a testbed for exploring neuro-symbolic integration, where neural perception is paired with symbolic reasoning to reconcile fine-grained visual cues with globally coherent program execution.
-In sum, CrochetBench and CrochetPARADE highlight a pronounced frontier at the intersection of multimodal learning, program synthesis, and digital fabrication.
-Progress here may one day close the loop from human intent to automated textile production, advancing both structured multimodal reasoning and computational craft.
-Ethics Statement
-We acknowledge that the original crochet pattern PDFs are protected under copyright and therefore do not distribute raw files or full texts.
-Instead, we release only structured JSON annotations generated with GPT, reference URLs to the original sources, and our parsing and annotation scripts.
-The benchmark is provided strictly for non-commercial academic use.
-This approach enables reproducible research while respecting intellectual property and ensuring that our dataset serves as a tool for studying structured generation rather than redistributing creative works.
-Reproducibility Statement
-We have taken several steps to ensure the reproducibility of our results.
-All datasets, task templates, and evaluation procedures are documented in the main text and appendix.
-An anonymous repository containing the full source code, experiment scripts, and detailed reproduction instructions has been made publicly available at: https://github.com/Peiyu-Georgia-Li/crochetBench.
-This ensures that all reported results can be independently verified and extended by the research community.
-Appendix A Related Work
-A.1 Multi-modal Datasets Beyond Captioning
-Most multimodal benchmarks have centered on descriptive pairing of images and natural language.
-Large-scale resources such as COCO [lin2015microsoftcococommonobjects] and Flickr30k [plummer2016flickr30kentitiescollectingregiontophrase] provide dense captions of everyday scenes, advancing vision–language representation learning.
-More recent datasets extend beyond captioning to procedural or instructional domains.
-Recipe1M+ aligns food images with ingredient lists and cooking steps [marin2019recipe1mdatasetlearningcrossmodal], while instructional video corpora such as YouCook2 [zhou2017automaticlearningproceduresweb] and HowTo100M [miech2019howto100mlearningtextvideoembedding] pair narrated demonstrations with visual segments.
-These resources emphasize semantic alignment but generally evaluate with retrieval- or similarity-based metrics.
-Our benchmark departs from this paradigm by pairing images with executable procedures.
-Rather than asking models to generate a semantically similar description, we require them to synthesize a program (CrochetPARADE DSL) that can be rendered and structurally verified.
-This shift enables functional evaluation—akin to program synthesis—and reduces reliance on subjective similarity measures.
-A.2 Bridging Visual Reasoning and Procedural Language
-Procedural understanding benchmarks highlight the importance of sequential, state-dependent reasoning.
-Datasets such as Assembly101 [sener2022assembly101largescalemultiviewvideo] and IKEA-ASM [benshabat2023ikeaasmdatasetunderstanding] capture human assembly activities, modeling dependencies across actions, objects, and preconditions.
-Instructional video benchmarks further test long-horizon understanding and error detection [wei2025instructionbenchinstructionalvideounderstanding, lin2024videoguibenchmarkguiautomation].
-Our task complements this line of work by grounding supervision not in temporally segmented actions but in artifact-centric procedures—finished crochet items paired with symbolic, stepwise instructions.
-This enables models to reason about topology, geometry, and sequential dependencies in a single unified representation.
-A.3 Crafting and Domain-Specific Procedural Data
-Closer to our domain, prior work has begun to explore fiber crafts.
-Seitz et al.
-introduced Digital Crochet, a visual, graph-based notation system for representing crochet patterns [digital_crochet].
-In knitting, Kaspar et al.
-developed Neural Inverse Knitting, mapping images of knitted patterns to machine instructions [kaspar2019neuralinverseknittingimages].
-These works underscore the feasibility of executable supervision in crafts, but remain limited in scale and scope.
-Our benchmark builds on this foundation by providing thousands of real, community-tested crochet patterns, along with a compiler and renderer for executable evaluation.
-A.4 Relation to Code Benchmarks
-The executable nature of CrochetPARADE connects it directly to program synthesis benchmarks.
-HumanEval [chen2021evaluatinglargelanguagemodels], MBPP [austin2021programsynthesislargelanguage], and APPS [hendrycks2021measuringcodingchallengecompetence] evaluate code generation by execution against unit tests.
-Similarly, Spider [yu2019spiderlargescalehumanlabeleddataset] and CoNaLa [yin2018mining] frame natural language to code translation tasks.
-In vision, Im2LaTeX-100K [deng2017imagetomarkupgenerationcoarsetofineattention] and pix2code [beltramelli2017pix2codegeneratingcodegraphical] evaluate image-to-program translation with render fidelity as the metric.
-Our tasks—ranging from text-to-DSL translation to image-to-stitch recognition—extend this paradigm to the crafting domain, introducing structured 2D/3D topology as part of the evaluation.
-This grounds multimodal learning in a setting where success requires both semantic alignment and structural correctness.
-A.5 Domain-Specific Languages for Executable Evaluation
-A key enabler of our benchmark is the use of a domain-specific language.
-Prior DSL-based benchmarks, such as Im2LaTeX and the CROHME competition on handwritten math recognition [crohme], demonstrate how symbolic formalisms enable reproducible rendering and structural evaluation.
-CrochetPARADE adopts this principle for fiber crafts: each pattern compiles into an abstract syntax tree that can be rendered and tested, supporting metrics such as ExecPass@K, structural unit tests, and image–render similarity.
-This functional perspective moves beyond surface-level similarity to test whether a model’s output actually works, in contrast to chemistry and recipe domain, which require quite amount of time to validate it [huang2024application, chefFusion, huang2025chemhgnn].
-Appendix B Additional Dataset Statistics
+Beyond raw instructions, the structured schema also records rich metadata.
+A representative dataset entry is provided in Table 4.
 B.1 Instruction Complexity by Skill Level
+Table 3 shows the instruction complexity by skill level.
 Skill Level Avg. Length Median Length Avg. Abbr. Count
@@
 Table 3: Instruction complexity by skill level. Length is measured in characters.
-Example Dataset Entry
-Table 10: Representative pattern entry from CrochetBench.
+Figure 6: Distribution of the top-10 most common project types in CrochetBench.
+Figure 7: Skill level distribution across the CrochetBench dataset.
+B.2 Example Dataset Entry
+Table 4 demonstrate a crochet dataset example.
 Field Value
@@
 Source input_file/Bags+Purses/SCC0303-005314M.pdf
-Instructions (truncated for brevity)
-B.2 Skill Level Distribution
-B.2.1 Overall Distribution
+Instructions Instructions: Note: Ch 2 at beg of each rnd counts as hdc.
+BAG With MC, ch 4.
+Join with sl st to form ring.
+1st rnd: Ch 2.
+11 hdc in ring.
+Join with sl st to top of ch 2.
+12 hdc.
+2nd rnd: Ch 2.
+1 hdc in same sp as sl st.
+2 hdc in each hdc around.
+Join.
+24 hdc.
+3rd rnd: Ch 2.
+1 hdc in same sp.
+1 hdc in next hdc.
+*2 hdc in next hdc, 1 hdc in next.* Rep around.
+Join.
+36 hdc.
+4th rnd: Ch 2.
+1 hdc in each hdc around.
+Join.
+5th rnd: Ch 2.
+1 hdc in same sp.
+1 hdc in next 2 hdc.
+*2 hdc, 1 hdc in next 2.* Join.
+48 hdc.
+6th rnd: As 4th rnd.
+7th rnd: Ch 2.
+1 hdc in next 2 hdc.
+*2 hdc, 1 hdc in next 3.* Rep.
+60 hdc.
+8th rnd: Ch 2.
+Back loops only, 1 hdc around.
+Join.
+9th–13th rnds: Ch 2.
+1 hdc in each hdc around.
+Join.
+14th rnd: Ch 2.
+1 hdc in same sp.
+1 hdc in next 4 hdc.
+*2 hdc, 1 hdc in next 4.* Join.
+72 hdc.
+15th rnd: Ch 2.
+1 hdc in same sp.
+1 hdc in next 5 hdc.
+*2 hdc, 1 hdc in next 5.* Join.
+84 hdc.
+16th–22nd rnds: Ch 2.
+1 hdc in each hdc around.
+Join.
+23rd rnd: Ch 2.
+1 hdc in next 4 hdc.
+*Hdc2tog, 1 hdc in next 5.* Rep.
+Hdc2tog.
+Join.
+72 sts.
+24th rnd: Ch 2.
+1 hdc in next 3 hdc.
+*Hdc2tog, 1 hdc in next 4.* Rep.
+Join.
+60 sts.
+25th rnd: Ch 2.
+1 hdc in next 2 hdc.
+*Hdc2tog, 1 hdc in next 3.* Rep.
+Join.
+48 sts.
+Fasten off.
+Eyes (Make 2) With A, ch 8.
+1st rnd: 2 sc in 2nd ch from hook.
+1 sc in next 5 ch.
+3 sc in last ch.
+Continue on rem loops, 1 sc in each ch.
+Join.
+17 sc.
+2nd rnd: Ch 1.
+3 sc in first sc.
+1 sc in next 7 sc.
+3 sc in next sc.
+1 sc in next 8 sc.
+Join.
+Fasten off.
+Handle With MC, ch 45.
+1st row: 1 sc in 2nd ch from hook.
+1 sc across.
+44 sc.
+Turn.
+2nd row: Ch 1.
+1 sc across.
+Turn.
+Rep last row 4 more times.
+Fasten off.
+Finishing Sew Eyes to Bag.
+Embroider mouth and teeth with A.
+Attach Handle.
+Cut cardboard circle to fit bottom.
+Table 4: Representative pattern entry from CrochetBench.
+B.3 Skill Level Distribution
+B.3.1 Overall Distribution
 Table 5 summarizes the overall distribution of skill levels across the CrochetBench dataset.
 The majority of patterns are labeled as easy (58.7%), followed by intermediate (32.3%).
-Only a small fraction are classified as beginner (7.6%) or experienced (1.3%).
-[^3] Three additional rare labels were observed: easy to intermediate (1 pattern), beginners (1 pattern), and beginner/easy (1 pattern).
+Only a small fraction are classified as beginner (7.6%) or experienced (1.3%).***Three additional rare labels were observed: easy to intermediate (1 pattern), beginners (1 pattern), and beginner/easy (1 pattern).
 Together they account for $0.1\%$ of the dataset.
-Table 11: Overall skill level distribution. Percentages are relative to all patterns with annotated skill levels.
 Skill Level Count Percentage
@@
 Total 6084 100%
+Table 5: Overall skill level distribution. Percentages are relative to all patterns with annotated skill levels.
 One pattern (0.02%) is missing an annotated skill level.
-B.2.2 Distribution by Project Type
+B.3.2 Distribution by Project Type
 We further break down skill levels by the top 10 most common project types.
@@
 Intermediate is the second most common, while beginner and experienced remain consistently low across categories.
-Table 12: Skill level distribution by top 10 project types. Percentages are within each project category.
 Project Type Easy Intermediate Beginner Experienced
@@
 Dishcloths 62.6% 27.5% 9.9%
+Table 6: Skill level distribution by top 10 project types. Percentages are within each project category.
 Overall, the predominance of easy patterns reflects the accessibility of crochet as a craft and aligns with the goal of many project types to cater to a wide audience.
 The relative scarcity of experienced-level patterns suggests that most published resources emphasize broad usability rather than advanced expertise.
-B.3 Pattern Complexity Analysis
-B.3.1 Instruction Length Statistics
+B.4 Pattern Complexity Analysis
+B.4.1 Instruction Length Statistics
 We first analyze the distribution of instruction lengths, measured in raw character counts.
@@
 The most complex patterns extend beyond 30,000 characters, while some very short patterns are as small as 20 characters.
-Table 13: Instruction length statistics (in characters).
+Out of 6,085 total patterns, 6,084 (99.98%) contain full instructions.
 Statistic Value
@@
 90th percentile 6403.9
-Out of 6,085 total patterns, 6,084 (99.98%) contain full instructions.
-B.3.2 Abbreviation Statistics
+Table 7: Instruction length statistics (in characters).
+B.4.2 Abbreviation Statistics
 Abbreviations, such as sc, dc, and hdc, are a distinctive element of crochet instructions.
@@
 Most patterns contain about 10 abbreviations, with values ranging from 1 to 31.
-Table 14: Abbreviation count statistics.
 Statistic Value
@@
 Max 31
-B.3.3 Complexity by Skill Level
+Table 8: Abbreviation count statistics.
+B.4.3 Complexity by Skill Level
 Instruction length correlates with the designated skill level.
@@
 Rare categories such as easy to intermediate skew extremely long due to outliers.
-Table 15: Instruction length and abbreviation counts by skill level.
 Skill Level Avg. Length Median Length Avg. Abbr. Count
@@
 Beginner/Easy 1063.0 1063.0 1
-B.3.4 Most and Least Complex Project Types
+Table 9: Instruction length and abbreviation counts by skill level.
+B.4.4 Most and Least Complex Project Types
 Finally, we identify the most complex and simplest project types by average instruction length.
-Tables 16 and 17 list the top 10 categories.
+Tables 10 and 11 list the top 10 categories.
 Garments such as dresses, vests, pants, and tunics are the most demanding, with average instructions exceeding 5,800 characters.
 By contrast, smaller accessories such as cowls, washcloths, scarves, and headbands are substantially shorter, typically under 2,000 characters.
-Table 16: Top 10 most complex project types (by average instruction length).
 Project Type Avg. Length Median Count
@@
 Aprons 4467.8 4494.0 11
-Table 17: Top 10 simplest project types (by average instruction length).
+Table 10: Top 10 most complex project types (by average instruction length).
 Project Type Avg. Length Median Count
@@
 Tech Accessories 2011.1 2099.0 13
+Table 11: Top 10 simplest project types (by average instruction length).
 Taken together, these results highlight strong alignment between project type, designated skill level, and instruction length.
@@
 (50 sc) **Finish:** Fasten off.
-DSL ¶ch.B ¶sc@B.A,5sc@B,ss@A ¶ch.A,sk,6sc2inc,ss@A ¶ch.A,sk,[sc2inc,sc]*6,ss@A ¶ch.A,sk,[sc2inc,2sc]*6,ss@A ¶ch.A,sk,24sc,ss@A ¶ch.A,sk,[sc2inc,3sc]*6,ss@A ¶[ch.A,sk,30sc,ss@A ¶]*2 ¶ch.A,sk,[scbl,scbl@[@],2scbl]*10,ss@A ¶ch.A,sk,40sc,ss@A ¶ch.A,sk,[sc2inc,3sc]*10,ss@A USER PROMPT Now generate DSL code for the following: [Image] [Instructions] Rnd 1: Ch 2, 6 sc in ring Rnd 2: 2 sc in each (12) Rnd 3: [Sc, sc, inc] around (16) Rnd 4: [Tr, sc] repeat around [DSL]
+DSL ¶ch.B ¶\$k=0\$,sc@B.A[k],5sc@B,ss@A[k++] ¶ch.A[k],sk,6sc2inc,ss@A[k++] ¶ch.A[k],sk,[sc2inc,sc]*6,ss@A[k++] ¶ch.A[k],sk,[sc2inc,2sc]*6,ss@A[k++] ¶ch.A[k],sk,24sc,ss@A[k++] ¶ch.A[k],sk,[sc2inc,3sc]*6,ss@A[k++] ¶[ch.A[k],sk,30sc,ss@A[k++] ¶]*2 ¶ch.A[k],sk,[scbl,scbl@[@],2scbl]*10,ss@A[k++] ¶ch.A[k],sk,40sc,ss@A[k++] ¶ch.A[k],sk,[sc2inc,3sc]*10,ss@A[k++] ¶ch.A[k],sk,[2sc,3sc]*10,ss@A[k++] USER PROMPT Now generate DSL code for the following: [Image] [Instructions] Rnd 1: Ch 2, 6 sc in ring Rnd 2: 2 sc in each (12) Rnd 3: [Sc, sc, inc] around (16) Rnd 4: [Tr, sc] repeat around [DSL]
 Appendix D CrochetPARADE: Pattern Renderer, Analyzer, and Debugger
-CrochetPARADE (short for Crochet Pattern Renderer, Analyzer, and Debugger) is an interactive platform that enables users to author, visualize, test, and export crochet patterns in both 2D and 3D [crochetparade].
-By combining a custom pattern grammar with simulation and rendering tools, CrochetPARADE addresses common issues of ambiguity, correctness, and interpretability in textual crochet instructions.
-[^4] https://www.crochetparade.org/
+Figure 8: Example of the CrochetBench translation pipeline.
+(Left) Natural language crochet instructions from the dataset.
+(Second) Automatically translated into CrochetPARADE DSL, a formal stitch grammar.
+(Third) Mesh rendering generated from the DSL.
+(Right) Target crocheted item image provided in the dataset.
+This pipeline enables direct text-to-image consistency checks, automated validation, and future training of NL $\rightarrow$ DSL models, analogous to text-to-code generation.
+CrochetPARADE (short for Crochet Pattern Renderer, Analyzer, and Debugger) is an interactive platform that enables users to author, visualize, test, and export crochet patterns in both 2D and 3D (Tassev, 2025).
+By combining a custom pattern grammar with simulation and rendering tools, CrochetPARADE addresses common issues of ambiguity, correctness, and interpretability in textual crochet instructions.***https://www.crochetparade.org/
 Core Capabilities.
@@
 This enables evaluation beyond surface-level metrics (e.g., BLEU, ROUGE) toward executor correctness—whether a generated pattern is valid, renderable, and stitch-balanced.
-Figure 3: Example of the CrochetBench translation pipeline.
-(Left) Natural language crochet instructions from the dataset.
-(Second) Automatically translated into CrochetPARADE DSL, a formal stitch grammar.
-(Third) Mesh rendering generated from the DSL.
-(Right) Target crocheted item image provided in the dataset.
-This pipeline enables direct text-to-image consistency checks, automated validation, and future training of NL $\rightarrow$ DSL models, analogous to text-to-code generation.
 Appendix E DSL Error Taxonomy
-To better understand failure cases in Task D, we extend the validator’s error analysis with detailed subcategories and examples. Errors fall into four major groups:
-1. Syntax Structure Errors
+To better understand failure cases in Task D, we extend the validator’s error analysis with detailed subcategories and examples.
 Unbalanced Brackets.
@@
 Examples Unbalanced brackets: (sc,hc5,sltr)infl)
-Multiple References Without Parentheses.
+Multiple References.
 Improper formatting of references.
@@
 (2ndrow):Ch1.(1scbl) ineachchtoendofrow.Turn
-2. Stitch Definition Errors
-Stitch Not Found.
+Undefined Stitch.
 Undefined stitch types not in the dictionary.
 Examples ch1, ch3, scfp, hdc_bar
-3. Labeling and Reference Errors
+Variable Naming Conflict.
+Conflicts between variable names and stitch names.
+Example Error: variable name matches stitch name. For example, \$ch=0\$ cannot be used since ’ch’ is a stitch name.
 Label Not Found.
@@
 Example Label not found: C
-Reusing Labels Incorrectly.
+Non-Adjacent Labels.
 Same label used for non-adjacent stitches.
 Example Cannot use same label over non-adjacent stitches. Consider using different labels.
-4. Structural and Formatting Issues
-Turning Errors.
+Turning Issue.
 Misplaced turning commands.
 Example Turning can happen only at the end of a row.
-Variable Naming Issues.
-Conflicts between variable names and stitch names.
-Example Error: variable name matches stitch name. For example, \$ch=0\$ cannot be used since ’ch’ is a stitch name.
 Runtime Errors.
@@
 Examples Cannot read properties of null (reading ’0’) Cannot use ’in’ operator to search for ’attach_id’ in NaN
+Multiplier Issue.
+Improper formatting in multiplier.
+Examples Error: Exception during pattern parsing: Multiplier set, but no stitch found: ch.B
+Appendix F Example Model-Generated Crochet Instructions
+This appendix provides the full verbatim crochet instructions generated by each model for the Task C (Instruction Generation) case study (Figure 2).
+We include the complete outputs without truncation or post-editing to preserve long-horizon procedural dependencies and enable transparent qualitative inspection.
+Because many failure modes in crochet pattern generation only emerge after multiple rounds, partial excerpts can obscure structural inconsistencies and topology errors.
+Presenting the full instructions allows readers to directly assess how each model handles stateful pattern progression, global geometry, and motif repetition beyond surface-level textual fluency.
+Ground Truth Notes: Ch 2 at beg of rnd does not count as hdc.
+Ch 3 at beg of rnd counts as dc.
+To change colors, work to last 2 loops on hook of first color.
+Draw new color through last 2 loops and proceed with new color.
+1.
+Ch 2 at beg of rnd does not count as hdc.
+Stripe Pat: With A, work 4 rnds.
+With B, work 4 rnds.
+These 8 rnds form Stripe Pat.
+Ch 4.
+Join with sl st to first ch to form ring.
+See diagram on page 3.
+1st rnd: Ch 3.
+13 dc in ring.
+Join with sl st to top of ch 3.
+14 dc.
+2nd rnd: Ch 3.
+1 dc in same sp as last sl st.
+2 dc in each dc around.
+Join with sl st to top of ch 3.
+28 dc.
+3rd rnd: Ch 3.
+4 dc in same sp as last sl st.
+*Dc3tog.
+5 dc in next dc.* Rep from * to last 3 sts.
+Dc3tog.
+Join with sl st to top of ch 3.
+4th rnd: Sl st in next dc.
+Ch 2.
+3 hdc in same dc as sl st.
+*Bobble in next st.
+3 hdc in next dc.
+Hdc3tog.* 3 hdc in next dc.
+Rep from * around, ending last rep at **.
+Join with sl st to first hdc.
+5th rnd: Sl st in next st.
+Ch 3.
+*3 dc in next st.
+1 dc in next st.
+3 dc in next st.
+1 dc in next st.
+Dc3tog.* 1 dc in next st.
+Rep from * around, ending last rep at **.
+Join with sl st to top of ch 3.
+6th rnd: Sl st in next st.
+Ch 2.
+1 hdc in same st as sl st.
+*(Bobble in next st.
+3 hdc in next st) twice.
+Bobble in next st.
+1 hdc in next st.
+Hdc3tog.* 1 hdc in next st.
+Rep from * around, ending last rep at **.
+Join with sl st to first hdc.
+7th rnd: Sl st in next st.
+Ch 3.
+1 dc in each of next 2 sts.
+*3 dc in next st.
+1 dc in next st.
+3 dc in next st.
+1 dc in each of next 3 sts.
+Dc3tog.* 1 dc in each of next 3 sts.
+Rep from * around, ending last rep at **.
+Join with sl st to top of ch 3.
+8th rnd: Sl st in next st.
+Ch 2.
+1 hdc in same st as sl st.
+*Bobble in next st.
+1 hdc in next st.
+(Bobble in next st.
+3 hdc in next st) twice.
+(Bobble in next st.
+1 hdc in next st) twice.
+Hdc3tog.* 1 hdc in next st.
+Rep from * around, ending last rep at **.
+9th rnd: Sl st in next st.
+Ch 3.
+1 dc in each of next 4 sts.
+3 dc in next st.
+1 dc in next st.
+PM in dc.
+3 dc in next st.
+1 dc in each of next 5 sts.
+Dc3tog.
+PM in dc3tog.
+*1 dc in each of next 5 sts.
+3 dc in next st.
+1 dc in next st.
+PM in dc.
+3 dc in next st.
+1 dc in each of next 5 sts.
+Dc3tog.
+PM in dc3tog.* Rep from * around.
+Join with sl st to top of ch 3.
+10th rnd: Sl st in next st.
+*1 hdc in next st.
+Bobble in next st.* Rep from * to * until 1 st before next marked st.
+3 hdc in next st.
+Bobble in next st.
+PM in Bobble.
+3 hdc in next st.
+*Bobble in next st.
+1 hdc in next st.* Rep from ** to 1 st before next marked st.
+Hdc3tog.
+PM in hdc3tog.
+Rep from * around.
+Join with sl st to first hdc.
+11th rnd: Sl st in next st.
+Ch 3.
+*1 dc in each st to 1 st before marked st.
+3 dc in next st.
+1 dc in next st.
+PM in dc.
+3 dc in next st.
+1 dc in each st until 1 st before next marked st.
+Dc3tog.
+PM in dc3tog.* Rep from * around.
+Join with sl st to top of ch 3.
+Rep 10th and 11th rnds, keeping cont of Stripe Pat, until work from center to outer point measures approx 30” [76 cm].
+Fasten off.
+Tassels (make 7): Cut a piece of cardboard 5” [12.5 cm] wide.
+Wind B around cardboard 40 times.
+Break yarn leaving a long end and thread end through needle.
+Slip needle through all loops and tie tightly.
+Remove cardboard and wind yarn tightly around loops 1” [2.5 cm] below fold.
+Fasten securely.
+Cut through remaining loops and trim ends evenly.
+Sew one tassel to each “peak” along the edge of Blanket as pictured.
+Gemini Special Stitch Bobble: [Yarn over, insert hook in indicated st, yarn over and pull up a loop, yarn over and draw through 2 loops] 5 times in same st, yarn over and draw through all 6 loops on hook.
+Instructions With Color A (Blue), make a magic ring.
+Rnd 1: Ch 1, 14 sc in ring.
+Join with sl st to first sc.
+[14 sc] Rnd 2: Ch 1, sc in first st.
+*Ch 2, skip 1 st, sc in next st.
+Rep from * around to last st, ch 2, skip last st.
+Join with sl st to first sc.
+Fasten off A.
+[7 sc, 7 ch-2 sps] Join Color B (Brown) with sl st in any ch-2 sp.
+Rnd 3: Ch 1, (Bobble, ch 2, Bobble) in same sp.
+*Sc in next sc, (Bobble, ch 2, Bobble) in next ch-2 sp.
+Rep from * around.
+Join with sl st to top of first Bobble.
+[14 Bobbles, 7 sc] Rnd 4: Sl st into ch-2 sp.
+Ch 1, (Bobble, ch 2, Bobble) in same sp.
+*Sc in next Bobble, Bobble in next sc, sc in next Bobble.
+(Bobble, ch 2, Bobble) in next ch-2 sp.
+Rep from * around, omitting last point group on final rep.
+Join with sl st to top of first Bobble.
+Fasten off B.
+[3 sts on each of 7 sides] Join Color A with sl st in any ch-2 sp.
+Rnd 5: Ch 1, (Bobble, ch 2, Bobble) in same sp.
+*(Sc in next st, Bobble in next st) twice, sc in next st.
+(Bobble, ch 2, Bobble) in next ch-2 sp.
+Rep from * around, omitting last point group on final rep.
+Join with sl st to top of first Bobble.
+[5 sts on each of 7 sides] Rnd 6: Sl st into ch-2 sp.
+Ch 1, (Bobble, ch 2, Bobble) in same sp.
+*(Sc in next st, Bobble in next st) three times, sc in next st.
+(Bobble, ch 2, Bobble) in next ch-2 sp.
+Rep from * around, omitting last point group on final rep.
+Join with sl st to top of first Bobble.
+Fasten off A.
+[7 sts on each of 7 sides] Join Color B with sl st in any ch-2 sp.
+Rnd 7: Ch 1, (Bobble, ch 2, Bobble) in same sp.
+*(Sc in next st, Bobble in next st) four times, sc in next st.
+(Bobble, ch 2, Bobble) in next ch-2 sp.
+Rep from * around, omitting last point group on final rep.
+Join with sl st to top of first Bobble.
+[9 sts on each of 7 sides] Rnd 8: Sl st into ch-2 sp.
+Ch 1, (Bobble, ch 2, Bobble) in same sp.
+*(Sc in next st, Bobble in next st) five times, sc in next st.
+(Bobble, ch 2, Bobble) in next ch-2 sp.
+Rep from * around, omitting last point group on final rep.
+Join with sl st to top of first Bobble.
+Fasten off B.
+[11 sts on each of 7 sides] Join Color A with sl st in any ch-2 sp.
+Rnd 9: Ch 1, (Bobble, ch 2, Bobble) in same sp.
+*(Sc in next st, Bobble in next st) six times, sc in next st.
+(Bobble, ch 2, Bobble) in next ch-2 sp.
+Rep from * around, omitting last point group on final rep.
+Join with sl st to top of first Bobble.
+[13 sts on each of 7 sides] Rnd 10: Sl st into ch-2 sp.
+Ch 1, (Bobble, ch 2, Bobble) in same sp.
+*(Sc in next st, Bobble in next st) seven times, sc in next st.
+(Bobble, ch 2, Bobble) in next ch-2 sp.
+Rep from * around, omitting last point group on final rep.
+Join with sl st to top of first Bobble.
+Fasten off.
+[15 sts on each of 7 sides] Finishing Weave in ends.
+Make 7 tassels with Color B and attach one to each ch-2 point of the star.
+GPT-4o Star-Shaped Crochet Blanket Pattern Materials: Yarn: Brown and Blue Crochet hook: Size H/8 (5.00 mm) Yarn needle Scissors Optional: Tassels (for points) Abbreviations: ch = chain sc = single crochet hdc = half double crochet dc = double crochet tr = treble sl st = slip stitch rep = repeat Instructions Foundation Ring: With Brown yarn, make a magic ring.
+Rnd 1: Ch 3 (counts as dc), 11 dc in ring.
+Join with sl st to top of ch-3.
+(12 dc) Rnd 2: Ch 3, dc in same space, 2 dc in each st around.
+Join.
+(24 dc) Rnd 3: Ch 3, *dc in next st, 2 dc in next st; rep from * around.
+Join.
+(36 dc) Rnd 4: Ch 3, *dc in next 2 sts, 2 dc in next st; rep from * around.
+Join.
+(48 dc) Change to Blue Yarn.
+Rnd 5: Ch 3, *dc in next 3 sts, 2 dc in next st; rep from * around.
+Join.
+(60 dc) Rnd 6: Ch 3, *dc in next 4 sts, 2 dc in next st; rep from * around.
+Join.
+(72 dc) Rnd 7: Ch 3, *dc in next 5 sts, 2 dc in next st; rep from * around.
+Join.
+(84 dc) Change to Brown Yarn.
+Rnd 8: Ch 3, *dc in next 6 sts, 2 dc in next st; rep from * around.
+Join.
+(96 dc) Rnd 9–11: Repeat Rnds 5–7 (using Blue & Brown alternately).
+Finishing: Fasten off and weave in ends.
+Tassels: Cut 4 lengths of Brown yarn (about 8 inches each).
+Wrap yarn around fingers, tie off center, trim ends, and attach one tassel to each point of the star.
+Block if necessary.
+Enjoy your star-shaped blanket!
+Claude 8-Point Star Afghan Materials: Worsted weight yarn in blue and brown Hook: Size H/8 (5mm) Center: With blue, ch 4, join with sl st to form ring.
+Rnd 1: Ch 1, 8 sc in ring, join with sl st.
+(8 sc) Rnd 2: Ch 1, 2 sc in each sc around, join.
+(16 sc) Rnd 3: Ch 1, *sc in next sc, 2 sc in next sc*, rep around, join.
+(24 sc) Rnd 4: Ch 1, *sc in next 2 sc, 2 sc in next sc*, rep around, join.
+(32 sc) Rnd 5: Ch 1, *sc in next 3 sc, 2 sc in next sc*, rep around, join.
+(40 sc) Rnd 6: Ch 1, *sc in next 4 sc, 2 sc in next sc*, rep around, join.
+(48 sc) Star Points Formation: Rnd 7: Ch 1, *sc in next 5 sc, 2 sc in next sc*, rep around, join.
+(56 sc) Rnd 8: Ch 1, *sc in next 6 sc, 2 sc in next sc*, rep around, join.
+(64 sc) Rnd 9: Change to brown.
+Ch 1, sc in each sc around, join.
+Rnd 10: Ch 1, *sc in next 7 sc, 2 sc in next sc*, rep around, join.
+(72 sc) Rnd 11: Ch 1, *sc in next 8 sc, 2 sc in next sc*, rep around, join.
+(80 sc) Rnd 12: Ch 1, *sc in next 9 sc, 2 sc in next sc*, rep around, join.
+(88 sc) Rnd 13: Change to blue.
+Ch 1, sc in each sc around, join.
+Rnd 14: Ch 1, *sc in next 10 sc, 2 sc in next sc*, rep around, join.
+(96 sc) Rnd 15: Ch 1, *sc in next 11 sc, 2 sc in next sc*, rep around, join.
+(104 sc) Rnd 16: Ch 1, *sc in next 12 sc, 2 sc in next sc*, rep around, join.
+(112 sc) Rnd 17: Change to brown.
+Ch 1, sc in each sc around, join.
+Rnd 18: Ch 1, *sc in next 13 sc, 2 sc in next sc*, rep around, join.
+(120 sc) Rnd 19: Ch 1, *sc in next 14 sc, 2 sc in next sc*, rep around, join.
+(128 sc) Rnd 20: Change to blue.
+Ch 1, sc in each sc around, join.
+Continue pattern of color changes every 3–4 rounds, increasing evenly to maintain flat star shape until desired size.
+Edging: Final Rnd: With blue, ch 1, *sc in next 2 sc, ch 3, sl st in 3rd ch from hook (picot made)*, rep around, join.
+Tassels: Cut 40 strands of brown yarn, 4 inches long.
+Divide into 8 groups of 5 strands each.
+Attach one tassel to each star point.
+Qwen-VL-7B Star Motif Crochet Pattern Materials: Worsted weight yarn in navy blue and brown Crochet hook size H (5.0 mm) or size needed to obtain gauge Yarn needle for weaving in ends Tassels (optional) Gauge: 16 sc and 16 hdc around a 4” circle Instructions: Round 1: Ch 10, join with sl st to form a ring.
+10 sc in the ring.
+Round 2: Ch 1, 1 sc in each sc around.
+Fasten off.
+Round 3: Ch 1, 1 hdc in each sc around.
+Fasten off.
+Round 4: Ch 1, 1 dc in each hdc around.
+Fasten off.
+Round 5: Ch 1, 1 tr in each dc around.
+Fasten off.
+Round 6: Ch 1, 1 hdc in each tr around.
+Fasten off.
+Round 7: Ch 1, 1 sc in each hdc around.
+Fasten off.
+Round 8: Ch 1, 1 hdc in each sc around.
+Fasten off.
+Round 9: Ch 1, 1 dc in each hdc around.
+Fasten off.
+Round 10: Ch 1, 1 tr in each dc around.
+Fasten off.
+Round 11: Ch 1, 1 hdc in each tr around.
+Fasten off.
+Round 12: Ch 1, 1 sc in each hdc around.
+Fasten off.
+Finishing: Join the rounds with slip stitches to form a star shape.
+Weave in ends.
+Attach tassels to each point of the star (optional).
+Note: Adjust the number of rounds to achieve the desired size of the star.
+DeepSeek-VL ch 100, sc in 2nd ch from hook, sc in each ch across, join with sl st in 1st sc, 100 loops made.
+Qwen-VL-72B Star Blanket Crochet Pattern Materials: Yarn in two colors: Blue (A) and Brown (B) Crochet hook size H/8 (5mm) Abbreviations: ch = chain sc = single crochet hdc = half double crochet dc = double crochet sl st = slip stitch Pattern Foundation Round: With Blue (A), ch 4, join with sl st to form a ring.
+Round 1: Ch 3 (counts as first dc), work 2 dc in ring, *ch 1, skip next ch, work 3 dc in ring; rep from * around, join with sl st to top of beg ch-3.
+(6 clusters) Round 2: Ch 3 (counts as first dc), work 2 dc in same st, *ch 1, skip next ch, work 3 dc in next cluster; rep from * around, join with sl st to top of beg ch-3.
+(12 clusters) Round 3: Ch 3 (counts as first dc), work 2 dc in same st, *ch 1, skip next ch, work 3 dc in next cluster; rep from * around, join with sl st to top of beg ch-3.
+(24 clusters) Round 4: Ch 3 (counts as first dc), work 2 dc in same st, *ch 1, skip next ch, work 3 dc in next cluster; rep from * around, join with sl st to top of beg ch-3.
+(48 clusters) Round 5: Ch 3 (counts as first dc), work 2 dc in same st, *ch 1, skip next ch, work 3 dc in next cluster; rep from * around, join with sl st to top of beg ch-3.
+(96 clusters) Color Change: Join Brown (B) with sl st in next ch-1 sp.
+Round 6: Ch 3 (counts as first dc), work 2 dc in same st, *ch 1, skip next ch, work 3 dc in next cluster; rep from * around, join with sl st to top of beg ch-3.
+(192 clusters) Round 7: Ch 3 (counts as first dc), work 2